### v1

# TransDF: Time-Series Forecasting Needs Transformed Label Alignment

Abstract
Training time-series forecasting models presents unique challenges in designing effective learning objectives. Existing methods predominantly utilize the temporal mean squared error, which faces two critical challenges: (1) label autocorrelation, which leads to bias from the label sequence likelihood; (2) excessive amount of tasks, which increases with the forecast horizon and complicates optimization. To address these challenges, we propose Transform-enhanced Direct Forecast (TransDF), which transforms the label sequence into decorrelated components with discriminated significance. Models are trained to align the most significant components, thereby effectively mitigating label autocorrelation and reducing task amount. Extensive experiments demonstrate that TransDF achieves state-of-the-art performance and is compatible with various forecasting models. Code is available at https://anonymous.4open.science/r/TransDF-88CF.
1 Introduction
Time-series forecasting involves predicting future data from historical observations and has been applied across diverse domains, such as weather forecasting in meteorology (application_weather, ), traffic flow prediction in transportation (application_traffic, ), and process monitoring in manufacturing (lsptd, ; spoti, ). To build effective forecast models, two questions warrant investigation: (1) How to design a neural network architecture to encode historical observations, and (2) How to devise a learning objective to train the neural network. Both are critical for model performance.
Recent research has primarily focused on developing neural network architectures. The key challenge lies in exploiting the autocorrelation in the historical sequences. To this end, various architectures have been proposed (qiu2024tfb, ; li2024foundts, ), such as recurrent neural networks (S4, ), convolutional neural networks (Timesnet, ; micn, ), and graph neural networks (FourierGNN, ). The current progress is marked by a debate between Transformers and simple linear models. Transformers, equipped with self-attention mechanisms, offer superior scalability (itransformer, ; PatchTST, ; fredformer, ). In contrast, linear models, which encapsulate temporal dynamics using linear layers, are straightforward to implement and often demonstrate strong performance (DLinear, ). These advancements showcase the rapid evolution in neural architecture design for time-series forecasting.
In contrast, the design of learning objectives has received less attention. Most existing methods follow the standard direct forecast (DF) paradigm, using the temporal mean squared error (TMSE) between forecast and label sequences as the learning objective (itransformer, ; PatchTST, ). While effective for many tasks, this approach has two critical issues in the context of time-series forecasting. First, the TMSE objective is biased against the true label sequence likelihood due to the presence of autocorrelation in the label sequence. Second, the number of prediction tasks increases with the forecast horizon, which complicates the optimization process since multitask learning is known to be challenging given excessive tasks (mtlsurvey, ; cagrad, ). These challenges present unique challenges in designing learning objectives for time-series forecasting.
To handle these challenges, we propose the Transform-enhanced Direct Forecast (TransDF), a simple yet effective refinement of the direct forecast paradigm. The key idea is to transform the label sequence into decorrelated components ranked by significance. By aligning the most significant decorrelated components, TransDF mitigates label autocorrelation and reduces the number of tasks, while retaining the efficiency and implementation simplicity of DF.
Our main contributions are summarized as follows:
  • We formulate two critical challenges in designing objectives for time-series forecasting: label autocorrelation that induces bias, and the excessive number of tasks that impedes optimization.
  • We propose TransDF, which transforms labels into decorrelated components with discriminated significance. By aligning the significant components, it addresses the two challenges above.
  • We conduct comprehensive experiments to demonstrate TransDF’s efficacy, consistently boosting the performance of state-of-the-art forecasting models across diverse datasets.
2 Preliminaries
This paper focuses on the time-series forecasting problem. By the way of preface, uppercase letters (e.g., $Y$) denote random variables, and bolded letters (e.g., $\mathbf{Y}$) denote matrices containing data or parameters. One key distinction warrants emphasis: we are concentrating on the design of learning objectives for training forecasting models fredf ; le2020probabilistic ; le2019shape , rather than on the design of neural network architectures to implement the forecast models itransformer ; DLinear .
Suppose $X$ is a time-series dataset with $\mathrm{D}$ covariates, where $X_{n}$ denotes the observation at the $n$-th step. At an arbitrary $n$-th step, the historical sequence is defined as $L=[X_{n-\mathrm{H}+1},\ldots,X_{n}]\in\mathbb{R}^{\mathrm{H}\times\mathrm{D}}$, the label sequence is defined as $Y=[X_{n+1},\ldots,X_{n+\mathrm{T}}]\in\mathbb{R}^{\mathrm{T}\times\mathrm{D}}$, where $\mathrm{H}$ is the historical length and $\mathrm{T}$ is the forecast horizon. The target of time-series forecasting is to train a model $g:\mathbb{R}^{\mathrm{H}\times\mathrm{D}}\rightarrow\mathbb{R}^{\mathrm{T}\times\mathrm{D}}$ that generates accurate prediction sequence $\hat{Y}$ approximating the label sequence.
There are two aspects to building forecast models: (1) neural network architectures that effectively encode historical sequences, and (2) learning objectives for training these neural networks. While this paper focuses on the learning objective, we provide a brief review of both aspects for contextualization.
2.1 Model architectures for time-series forecasting
Neural networks have been pervasive in encoding historical sequences for their capability of automating feature interactions and capturing nonlinear correlations. Notable examples include RNNs (e.g., S4 (S4, ), Mamba), CNNs (e.g., TimesNet (Timesnet, )), and GNNs (e.g., MTGNN (Mtgnn, )), each tailored to encode the dynamics within input sequences. The current progress centers on the comparison between Transformer-based and MLP-based architectures. Transformers (e.g., PatchTST (PatchTST, ), iTransformer (itransformer, )) exhibit substantial scalability with increasing data size but entail high computational costs. In contrast, MLPs (e.g., DLinear (DLinear, ), TimeMixer (wang2024timemixer, )) are generally more efficient but less scalable with larger datasets and struggle to handle varying input lengths.
2.2 Learning objectives for time-series forecasting
Modern time-series models predominantly adopt the direct forecast paradigm, generating T-step forecasts simultaneously using a multi-output head Informer ; DLinear ; itransformer . The learning objective is typically the temporal mean squared error (TMSE) between the forecast and label sequences, given by:
$\mathcal{L}_{\mathrm{tmp}}=\sum_{t=1}^{\mathrm{T}}\left(Y_{t}-\hat{Y}_{t}\right)^{2},$ (1)
which is widely employed in recent studies (e.g., FreTS (FreTS, ), iTransformer (itransformer, ), FredFormer (fredformer, )). However, this objective has been shown to be biased due to the autocorrelation present in the label sequence fredf . To address this bias, one line of research advocates for shape alignment between the forecast and label sequences to exploit autocorrelation (e.g., Dilate (le2019shape, ), Soft-DTW (soft-dtw, ) and DPTA (dtw, )). However, these methods lack rigorous theoretical guarantees for unbiased objective and empirical evidence of improved performance. Another notable approach involves computing the forecast error in the frequency domain, which reduces bias with theoretical guarantees (fredf, ).
3 Methodology
3.1 Motivation
The learning objective is a fundamental component in training effective forecasting models, yet its importance remains underexplored. Existing approaches predominantly employ the TMSE in (1) as the objective (itransformer, ; PatchTST, ). This practice, however, encounters two fundamental limitations rooted in the characteristics of time-series forecasting task.
First, TMSE introduces bias due to autocorrelation. In time-series forecasting, observations exhibit strong dependencies on their past values (DLinear, ), resulting in step-wise correlation in the label sequence. In contrast, TMSE treats the forecast of each step as an independent task, thereby neglecting these correlations. This mismatch makes TMSE biased with respect to the true likelihood of the label sequence, as presented in Theorem 3.1.
Theorem 3.1 (Autocorrelation bias).
Given label sequence $Y$ where $\Sigma\in\mathbb{R}^{\mathrm{T}\times\mathrm{T}}$ denotes the step-wise correlation coefficient, the TMSE in (1) is biased compared to the negative log-likelihood of the label sequence, which is given by:
$\mathrm{Bias}=\left\|Y-\hat{Y}\right\|_{\Sigma^{-1}}^{2}-\left\|Y-\hat{Y}\right\|^{2}-\frac{1}{2}\log\left|\Sigma\right|.$ (2)
where $\|Y-\hat{Y}\|_{\Sigma^{-1}}^{2}=(Y-\hat{Y})^{\top}\Sigma^{-1}(Y-\hat{Y})$. The bias vanishes if different steps in $Y$ are decorrelated. [^1] The pioneering work (fredf, ) identifies the bias under the first-order Markov assumption on the label sequence. This study generalizes this bias without the first-order Markov assumption.
Second, TMSE poses optimization difficulties as the forecast horizon grows. The large forecast horizon is crucial for applications such as manufacturing (enabling comprehensive production planning (lsptd, )) and transportation (enabling proactive traffic management (application_traffic, )). As TMSE treats each forecasted step as an independent task, a large horizon results in excessive tasks. However, optimization is known to be difficult given excessive tasks (mtlsurvey, ; cagrad, ), as gradients from different tasks often conflict (mtlsurvey, ; cagrad, ), impeding convergence and leading to suboptimal model performance.
Designing effective learning objectives to handle the two limitations is challenging. The recent pioneering work FreDF (fredf, ) proposes a frequency loss, which transforms the label and forecast sequences into frequency components and aligns them in the frequency domain. This approach is motivated by Theorem 3.1: bias vanishes if different components are decorrelated. However, the decorrelation of frequency components holds only when the forecast horizon $\mathrm{T}\rightarrow\infty$ (see Theorem 3.3 in (fredf, )). In real-world settings with finite horizon, frequency components remain correlated, rendering FreDF ineffective in eliminating bias. Additionally, the optimization difficulty remains, since transforming to the frequency domain retains the label length. Consequently, FreDF fails to address the autocorrelation bias and the optimization difficulty.
Given the critical role of objective in training forecast models and the limitations of existing approaches, it is compelling to develop an innovative objective to address the limitations and advance forecast performance. Importantly, there are two questions that warrant investigation. How to devise an objective that eliminates autocorrelation and reduces task amount? Does it improve forecast performance?
3.2 Transforming label sequence with optimized projection matrix
In this section, we present a method for transforming label sequences into latent components to eliminate autocorrelation and distinguish significant components. Suppose $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$ contains normalized label sequences for $\mathrm{m}$ samples, $\mathbf{P}=\left[\mathbf{P}_{1},\mathbf{P}_{2},...,\mathbf{P}_{\mathrm{T}}\right]$ is the projection matrix, and the components are produced as $\mathbf{Z}=\mathbf{Y}\mathbf{P}$. The target is for $\mathbf{Z}$ to be decorrelated and ranked by significance. For example, FreDF specifies $\mathbf{P}$ as a Fourier matrix (pswi, ), which does not adapt to specific data properties and thus fails to decorrelate the components and distinguish the significant components [^2] In the subsequent paragraphs, we use the univariate case with $\mathrm{D}=1$ for clarity. In the multivariate case, label sequences are concatenated along the forecast horizon to produce decorrelated components..
A natural approach to obtaining the projection matrix $\mathbf{P}$ is solving optimization problem with constraints to ensure the desired properties. To find the $p$-th component, the projection vector can be calculated by solving the following problem:
$\mathbf{P}_{p}^{*}=$ $\ \text{arg}\underset{\mathbf{P}_{p}}{\text{max}}$ $\left(\mathbf{Y}\mathbf{P}_{p}\right)^{\top}\left(\mathbf{Y}\mathbf{P}_{p}\right)$ (3)
  subject to $\begin{cases}\|\mathbf{P}_{p}\|^{2}=1&\\ \mathbf{P}_{p}^{\top}\mathbf{P}_{j}=0,\ \forall\,j<p&\text{if }p>1\end{cases}$
where $\mathbf{Z}_{p}=\mathbf{Y}\mathbf{P}_{p}$ is the $p$-th component, the normalization constraint $\|\mathbf{P}_{p}\|^{2}=1$ is imposed to avoid trivial solution: $\mathbf{P}_{p}\rightarrow\infty$. The optimization target is to maximize the variance of $\mathbf{Z}_{p}$, which is equivalent to maximizing its significance, as components with larger variance contain richer information. For $p>1$, the projection axis is required to be orthogonal to the previous axes to avoid redundancy. By solving the optimizations above from $p=1$ to $\mathrm{T}$ sequentially, we obtain the projection matrix $\mathbf{P}^{*}=\left[\mathbf{P}_{1}^{*},...,\mathbf{P}_{\mathrm{T}}^{*}\right]$. The components are then produced as $\mathbf{Z}=\mathbf{Y}\mathbf{P}^{*}$.
Lemma 3.2 (Decorrelated components).
Suppose $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$ contains normalized label sequences for $\mathrm{m}$ samples, $\mathbf{Z}=[\mathbf{Z}_{1},...,\mathbf{Z}_{\mathrm{T}}]$ are the obtained components; for any $p\neq p^{\prime}$, we have $\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=0$.
Lemma 3.3.
The projection matrix $\mathbf{P}^{*}$ can be obtained via singular value decomposition (SVD): $\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}(\mathbf{P}^{*})^{\top}$, where $\mathbf{U}\in\mathbb{R^{\mathrm{m\times m}}}$ and $\mathbf{P}^{*}\in\mathbb{R^{\mathrm{T\times T}}}$ consist of singular vectors, and the diagonal of $\mathbf{\Lambda}\in\mathbb{R^{\mathrm{m\times\mathrm{T}}}}$ consists of singular values.
Theoretical Justification.
According to Theorem 3.1, the bias vanishes as the correlations between labels to be aligned are eliminated. The obtained components are decorrelated (Lemma 3.2), thereby mitigating autocorrelation-induced bias. Moreover, component significance escalates from $\mathbf{Z}_{1}$ to $\mathbf{Z}_{\mathrm{T}}$ as they are derived by maximizing significance under sequentially augmented constraints. Furthermore, $\mathbf{P}^{*}$ can be computed via SVD (Lemma 3.3), offering an efficient alternative to sequentially solving the constrained optimization problems in (3).
Case study. To showcase the implications of the obtained components, a case study was conducted on the ETTh1 dataset. Implementation details are provided in Appendix A. The results are illustrated in LABEL:fig:auto, with key observations summarized as follows:
  • Decorrelation effect: LABEL:fig:auto (a) compares the correlation volume in the label sequence and the generated components. In the left panel, the value at row i and column j represents the correlation between the i-th and j-th steps. A large number of non-diagonal elements exhibit substantial values, with approximately 50.5% exceeding 0.25, indicating notable autocorrelation in the label sequence. In contrast, the right panel shows negligible values for the non-diagonal elements. This demonstrates that transforming the label sequence into components effectively eliminates correlation, thereby corroborating Theorem 3.2.
  • Significance Discrimination: LABEL:fig:auto (b) compares the variance of the label sequence and associated components. In the left panel, the variance of different steps in the label sequence is relatively uniform, ranging from -1.5 to 1.5, suggesting that all steps are equally significant. In the right panel, however, only a few components exhibit large variance, while the others fluctuate within a narrow range. This indicates that the significance of different components can be clearly discerned, allowing for a trade-off between a slight loss of information and reduced optimization complexity by focusing on the most significant components.
The transformation is highly inspired by principal component analysis (pca1, ; pca2, ). However, one key distinction warrants emphasis. Existing works dominantly employ principal component analysis on input features for obtaining informative representations (pca3, ; pca4, ), in contrast, we adapt it to label sequence, specifically aiming to reduce autocorrelation bias and simplify optimization for time-series forecasting. To our knowledge, this remains a technically innovative strategy.
Algorithm 1 The workflow of TransDF.
Input: $\hat{\mathbf{Y}}$: forecast sequences, $\mathbf{Y}$: label sequences.
Parameter: $\alpha$: the relative weight of the transformed loss, $\gamma$: the ratio of involved significant components.
Output: $\mathcal{L}_{\alpha,\gamma}$: the obtained learning objective.
1:  $\mathbf{Y}\leftarrow\mathrm{Standardize}(\mathbf{Y})$.
2:  $\mathbf{P}^{*}\leftarrow\mathrm{SVD(\mathbf{Y})}$
3:  $\mathbf{Z}\leftarrow\mathbf{Y}\mathbf{P}^{*},\hat{\mathbf{Z}}\leftarrow\hat{\mathbf{Y}}\mathbf{P}^{*}$
4:  $\mathrm{K}\leftarrow\mathrm{round}(\gamma\cdot\mathrm{T})$
5:  $\mathcal{L}_{\mathrm{trans},\gamma}\leftarrow\|\hat{\mathbf{Z}}_{\cdot,1:\mathrm{K}}-\mathbf{Z}_{\cdot,1:\mathrm{K}}\|_{1}$
6:  $\mathcal{L}_{\mathrm{tmp}}\leftarrow\|\hat{\mathbf{Y}}-\mathbf{Y}\|_{2}^{2}$
7:  $\mathcal{L}_{\alpha,\gamma}:=\alpha\cdot\mathcal{L}_{\mathrm{trans},\gamma}+(1-\alpha)\cdot\mathcal{L}_{\mathrm{tmp}}.$
3.3 Model implementation
In this section, we present the implementation details of TransDF. The approach centers on extracting the latent components from the label sequence, then optimizing the forecasting model using the most significant components.
Given an input historical sequence, the forecast model predicts a sequence $\hat{\mathbf{Y}}$. In line with prevailing preprocessing practices itransformer ; DLinear ; fredformer , label sequences are first standardized (step 1), which facilitates the decorrelation prerequisite specified in Lemma 3.2. Next, following Lemma 3.3, we compute the optimal projection by applying SVD to the label sequence. The matrix $\mathbf{P}^{*}$, composed of the right singular vectors, provides the required projections described in (3). Both forecasted and label sequences are then projected into the latent component space (step 3), where the first column carries the largest significance, successively diminishing across columns.
Suppose $\mathrm{K}$ is the number of retained components, the training objective using them is given by:
$\mathcal{L}_{\mathrm{trans},\gamma}:$ $=\left\|\hat{\mathbf{Z}}_{\cdot,1:\mathrm{K}}-\mathbf{Z}_{\cdot,1:\mathrm{K}}\right\|_{1},$ (4)
where $\mathrm{K}=\mathrm{round}(\gamma\cdot\mathrm{T})$, with $\gamma$ controlling the involution ratio, the $\ell_{1}$ norm $\left\|\cdot\right\|_{1}$ computes the sum of element-wise absolute differences. Typically, we use the $\ell_{1}$ norm instead of the squared norm following (fredf, ), considering that latent components typically vary greatly in scale (LABEL:fig:auto), which makes the squared norm unstable in practice. The $\ell_{1}$ norm yields more stable and robust optimization.
Finally, the objectives calculated with the transformed components and the label sequence are fused following (fredf, ), with $0\leq\alpha\leq 1$ controlling the relative contribution:
$\mathcal{L}_{\alpha,\gamma}:=\alpha\cdot\mathcal{L}_{\mathrm{trans},\gamma}+(1-\alpha)\cdot\mathcal{L}_{\mathrm{tmp}}.$ (5)
By projecting both forecasts and labels into a decorrelated latent space, TransDF effectively reduces autocorrelation bias. By focusing exclusively on the most significant components, TransDF reduces optimization difficulty with minimal information sacrifice. Key advantages of the canonical DF framework itransformer ; DLinear , like efficient inference and multi-task learning capabilities, are preserved. Crucially, TransDF is model-agnostic, offering practitioners the flexibility to employ the most suitable forecasting model for each specific scenario.
4 Experiments
To demonstrate the efficacy of TransDF, there are six aspects empirically investigated:
  1. 1.
    Performance: Does TransDF work? We compare TransDF with state-of-the-art baselines using public datasets on long-term forecasting in Section 4.2 and short-term forecasting tasks in Appendix E.2. Moreover, we compare it with other learning objectives in Section 4.3.
  2. 2.
    Gain: How does it work? Section 4.4 offers an ablative study to dissect the contributions of the individual factors of TransDF, elucidating their roles in enhancing forecast accuracy.
  3. 3.
    Generality: Does it support other forecasting models? Section 4.5 verifies the adaptability of TransDF across different forecasting models, with additional results documented in Appendix E.3.
  4. 4.
    Flexibility: Does it support alternative transformations? Section 4.5 also investigates generating latent components with other transformations to showcase flexibility of implementation.
  5. 5.
    Sensitivity: Does it require careful fine-tuning? Section 4.6 presents a sensitivity analysis of the hyperparameter $\alpha$, where TransDF maintains efficacy across a broad range of parameter values.
  6. 6.
    Efficiency: Is TransDF computationally expensive? Section E.1 investigates the running cost of TransDF in diverse settings.
4.1 Setup
Datasets.
In this work, we conduct experiments on public datasets following prior works Timesnet . For long-term forecasting, we use ETT (4 subsets), ECL, Traffic, Weather, and PEMS (itransformer, ). For short-term forecasting, we use M4. All datasets are split chronologically into training, validation, and testing sets. Detailed dataset statistics are provided in Appendix D.1.
Baselines.
We compare TransDF against several established methods, grouped as: (1) Transformer-based methods: Transformer (Transformer, ), Autoformer (Autoformer, ), FEDformer (fedformer, ), iTransformer (itransformer, ), and Fredformer (fredformer, ); (2) MLP-based methods: DLinear (DLinear, ), TiDE (das2023long, ), and FreTS (FreTS, ); and (3) other competitive models: TimesNet (Timesnet, ) and MICN (micn, ).
Implementation.
The baseline models are reproduced using the scripts provided by Fredformer (fredformer, ). Notably, we disable the drop-last trick to ensure fair comparison following Qiu et al. qiu2024tfb . They are trained using the Adam (Adam, ) optimizer to minimize the TMSE loss. Datasets are split chronologically into training, validation, and test sets. Following the protocol outlined in the comprehensive benchmark (qiu2024tfb, ), the dropping-last trick is disabled during the test phase. When integrating TransDF to enhance an established model, we adhere to the associated hyperparameter settings in the public benchmark (fredformer, ; itransformer, ), only tuning $\alpha$, $\gamma$ and learning rate conservatively. Experiments are conducted on Intel(R) Xeon(R) Platinum 8383C CPUs and NVIDIA RTX H100 GPUs. More implementation details are provided in Appendix D.2.
4.2 Overall performance
Table 1: Long-term forecasting performance.
Models
PDF
Fredformer
iTransformer
FreTS
TimesNet
MICN
TiDE
DLinear
FEDformer
Autoformer
Transformer
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
0.380
0.393
0.387 0.398
0.411
0.414
0.414
0.421
0.438
0.430
0.396
0.421
0.413
0.407
0.403
0.407
0.442
0.457
0.526
0.491
0.799
0.648
ETTm2
0.272
0.317
0.280 0.324
0.295
0.336
0.316
0.365
0.302
0.334
0.308
0.364
0.286
0.328
0.342
0.392
0.308
0.354
0.315
0.358
1.662
0.917
ETTh1
0.431
0.429
0.447 0.434
0.452
0.448
0.489
0.474
0.472
0.463
0.533
0.519
0.448
0.435
0.456
0.453
0.447
0.470
0.477
0.483
0.983
0.774
ETTh2
0.359
0.388
0.377
0.402
0.386
0.407
0.524
0.496
0.409
0.420
0.620
0.546
0.378
0.401
0.529
0.499
0.452
0.461
0.448
0.460
2.688
1.291
ECL
0.170
0.260
0.191
0.284
0.179 0.270
0.199
0.288
0.212
0.306
0.192
0.302
0.215
0.292
0.212
0.301
0.214
0.328
0.249
0.354
0.265
0.358
Traffic
0.419
0.280
0.486
0.336
0.426 0.285
0.538
0.330
0.631
0.338
0.529
0.312
0.624
0.373
0.625
0.384
0.640
0.398
0.662
0.416
0.692
0.379
Weather
0.241
0.280
0.261
0.282
0.269
0.289
0.249
0.293
0.271
0.295
0.264
0.321
0.272
0.291
0.265
0.317
0.326
0.372
0.319
0.365
0.699
0.601
PEMS03
0.097
0.208
0.146
0.260
0.122
0.233
0.149
0.261
0.126
0.230
0.106 0.223
0.316
0.370
0.216
0.322
0.152
0.275
0.411
0.475
0.122
0.226
PEMS08
0.141
0.237
0.171
0.271
0.149 0.247
0.174
0.275
0.152
0.243
0.153
0.258
0.318
0.378
0.249
0.332
0.226
0.312
0.422
0.456
0.240
0.261
  • Note: We fix the input length as 96 following (itransformer, ). Bold and underlined denote best and second-best results, respectively. Avg indicates average results over forecast horizons: T=96, 192, 336 and 720. TransDF employs the top-performing baseline on each dataset as its underlying forecasting model.
Table 1 presents the long-term forecasting results. TransDF consistently improves base model performance. For example, on ETTh1, it reduces Fredformer’s MSE by 0.016. Similar gains across other datasets further validate its effectiveness. These results suggest that modifying the learning objective can yield improvements comparable to, or even exceeding, those from architectural advancements. We attribute this to two key aspects of TransDF: its decorrelation effect for debiased training, and its discrimination on significant components, which simplifies optimization.
Showcases.
We visualize the forecast sequences and the generated components to showcase the improvements of TransDF in forecast quality. A snapshot on ETTm2 with historical window $\mathrm{H}=96$ and forecast horizon $\mathrm{T}=336$ is depicted in LABEL:fig:case. Although the model trained using canonical DF captures general trends, its forecast struggles with large variations (e.g., peaks within steps 100-400). This reflects its difficulty in modeling significant, high-variance components. In contrast, TransDF, by explicitly discriminating and aligning these significant components, generates a forecast that accurately captures these large variations, including the peaks within steps 100-400.
4.3 Learning objective comparison
Table 2: Comparable results with other objectives for time-series forecast.
Loss TransDF FreDF Koopman Dilate Soft-DTW DPTA DF
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Fredformer ETTm1 0.379 0.393 0.384 0.394 0.389 0.400 0.389 0.400 0.397 0.402 0.396 0.402 0.387 0.398
ETTh1 0.431 0.429 0.438 0.434 0.452 0.443 0.453 0.442 0.460 0.449 0.460 0.449 0.447 0.434
ECL 0.178 0.270 0.179 0.272 0.190 0.282 0.187 0.280 0.206 0.298 0.202 0.294 0.191 0.284
Weather 0.255 0.276 0.256 0.277 0.257 0.279 0.258 0.280 0.261 0.280 0.260 0.280 0.261 0.282
iTransformer ETTm1 0.395 0.401 0.405 0.405 0.413 0.416 0.407 0.412 0.417 0.415 0.416 0.415 0.411 0.414
ETTh1 0.438 0.434 0.442 0.437 0.455 0.451 0.452 0.448 0.470 0.457 0.463 0.454 0.452 0.448
ECL 0.170 0.260 0.176 0.264 0.178 0.269 0.178 0.269 0.175 0.266 0.177 0.267 0.179 0.270
Weather 0.251 0.272 0.257 0.276 0.289 0.313 0.286 0.309 0.292 0.316 0.291 0.313 0.269 0.289
  • Note: Bold and underlined denote best and second-best results, respectively. The reported results are averaged over forecast horizons: T=96, 192, 336 and 720.
Table 2 compares TransDF against other time-series learning objectives: FreDF (fredf, ), Koopman (koopman, ), Dilate (le2019shape, ), Soft-DTW (soft-dtw, ), and DPTA (dtw, ). For fair evaluation, we integrated their official implementations into both Fredformer and iTransformer.
Overall, shape alignment objectives (Dilate, Soft-DTW, DPTA) offer little performance gain over canonical DF (using TMSE loss), consistent with the findings in (le2019shape, ). This phenomenon is rationalized by the fact that they do not mitigate the label correlation nor reduce task amounts for simplifying optimization. FreDF improves performance by partly addressing autocorrelation bias. However, as discussed in Section 3.1, FreDF does not fully eliminate this bias, nor does it distinguish significant components to simplify the optimization landscape. TransDF directly addresses these two limitations of FreDF, leading to its superior overall performance.
4.4 Ablation studies
Table 3: Ablation study results.
Model Decorrelation Reduction Data T=96 T=192 T=336 T=720 Avg
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
DF ETTm1 0.326 0.361 0.365 0.382 0.396 0.404 0.459 0.444 0.387 0.398
ETTh1 0.377 0.396 0.437 0.425 0.486 0.449 0.488 0.467 0.447 0.434
ECL 0.150 0.242 0.168 0.259 0.182 0.274 0.214 0.304 0.179 0.270
Weather 0.174 0.228 0.213 0.266 0.270 0.316 0.337 0.362 0.249 0.293
TransDF ETTm1 0.338 0.366 0.369 0.383 0.397 0.403 0.458 0.441 0.391 0.398
ETTh1 0.376 0.395 0.437 0.430 0.478 0.450 0.469 0.467 0.440 0.436
ECL 0.150 0.239 0.164 0.253 0.178 0.268 0.210 0.296 0.175 0.264
Weather 0.170 0.216 0.213 0.259 0.262 0.300 0.332 0.351 0.244 0.281
TransDF ETTm1 0.324 0.359 0.362 0.379 0.390 0.400 0.451 0.438 0.382 0.394
ETTh1 0.373 0.395 0.433 0.423 0.476 0.445 0.474 0.463 0.439 0.431
ECL 0.147 0.238 0.162 0.252 0.174 0.267 0.205 0.294 0.172 0.263
Weather 0.172 0.220 0.211 0.259 0.261 0.301 0.331 0.353 0.244 0.283
TransDF ETTm1 0.321 0.357 0.360 0.378 0.389 0.400 0.447 0.435 0.379 0.393
ETTh1 0.368 0.391 0.424 0.422 0.467 0.441 0.465 0.463 0.431 0.429
ECL 0.145 0.235 0.159 0.249 0.173 0.264 0.203 0.292 0.170 0.260
Weather 0.169 0.219 0.210 0.258 0.259 0.297 0.327 0.349 0.241 0.280
  • Note: Bold and underlined denote best and second-best results, respectively.
Table 3 presents an ablation study dissecting the contributions of critical factors in TransDF: the decorrelation effect and the task reduction effect. The main findings are summarized as follows.
  • TransDF improves DF by reducing the number of tasks to optimize. To this end, it employs a randomized matrix as the projection matrix to generate components and aligns only a subset of the obtained components. The involution ratio $\gamma$ is finetuned on the validation set. It consistently improves over DF (e.g., --0.012 MAE on Weather). This demonstrates that reducing tasks with a minimal loss of label information can reduce optimization difficulty and improve performance.
  • TransDF improves DF by aligning decorrelated components. To this end, the objective is calculated in (5) with $\gamma=1$. It also outperforms DF, achieving the second-best results overall. This demonstrates aligning decorrelated label components to mitigate bias benefits forecast performance.
  • TransDF integrates both factors above by aligning the most significant decorrelated components, which achieves the best performance, demonstrating the synergistic effect of these two factors.
4.5 Generalization studies
In this section, we investigate the utility of TransDF with different transformation strategies and forecast models, to showcase the generality of TransDF. In the bar-plots, the forecast errors are averaged over forecast lengths (96, 192, 336, 720), with error bars as 50% confidence intervals.
Varying transformations.
Table 4: Varying transformations results.
ECL Weather
Trans MSE RI MAE RI MSE RI MAE RI
None 0.179 - 0.270 - 0.249 - 0.293 -
RPCA 0.171 4.31% 0.261 3.16% 0.244 1.78% 0.286 2.38%
SVD 0.175 2.24% 0.264 2.18% 0.248 0.34% 0.290 0.93%
FA 0.175 2.35% 0.265 1.82% 0.245 1.35% 0.287 1.97%
Ours 0.170 4.86% 0.260 3.57% 0.241 2.94% 0.280 4.28%
  • Note: RI refers to the relative improvement (error reduction) over the baseline. Bold and underlined denote best and second-best results.
We select alternative approaches to transform the label sequence into latent components and report the forecast performance in Table 4. The selected transformation methods include robust principal component analysis (RPCA) (rpca, ), SVD (svd, ), and factor analysis (FA, ). Noting that the output of SVD yields components here, not a projection matrix as in Section 3.2. Implementation details are in Appendix C. Overall, all these transformation methods outperform canonical DF without transformation. However, the components obtained by these methods, including RPCA, cannot be guaranteed to be decorrelated. Consequently, autocorrelation bias may persist. In contrast, our approach ensures full decorrelation of the derived components (see Lemma 3.2), effectively addressing autocorrelation bias and leading to the best overall performance.
Varying forecast models.
We explore the versatility of TransDF in augmenting representative forecast models: Fredformer (fredformer, ), iTransformer (itransformer, ), FreTS (FreTS, ), and DLinear (DLinear, ). As illustrated in LABEL:fig:backbone, TransDF improves forecast performance in all cases. For instance, on the Weather dataset, iTransformer and FreTS with TransDF achieve substantial reductions in MSE—up to 6.9% and 2.9%, respectively. Further evidence of TransDF’s versatility can be found in Appendix E.3. These results confirm TransDF’s potential as a plug-and-play strategy to enhance adiverse forecast models.
4.6 Hyperparameter sensitivity
Table 5: Hyperparameter results on $\alpha$.
ETTm1
ETTh2
Weather
$\alpha$
MSE MAE MSE MAE MSE MAE
0 0.3867 0.3979 0.3766 0.4019 0.2486 0.2930
0.3 0.3871 0.3983 0.3742 0.3982 0.2439 0.2851
0.5 0.3864 0.3976 0.3703 0.3964 0.2432 0.2833
0.7 0.3831 0.3959 0.3674 0.3943 0.2433 0.2849
1 0.3850 0.3933 0.3606 0.3890 0.2753 0.3209
  • Note: Bold and underlined denote best and second-best results.
Table 6: Hyperparameter results on $\gamma$.
ETTm1
ETTh2
Weather
$\gamma$
MSE MAE MSE MAE MSE MAE
0.1 0.3915 0.4002 0.3816 0.4029 0.2437 0.2845
0.3 0.3849 0.3964 0.3694 0.3961 0.2424 0.2825
0.5 0.3817 0.3943 0.3651 0.3923 0.2466 0.2877
0.7 0.3798 0.3930 0.3603 0.3886 0.2443 0.2861
1 0.3814 0.3940 0.3624 0.3903 0.2491 0.2924
  • Note: Bold and underlined denote best and second-best results.
In this section, we examine the impact of critical hyperparameters on the performance of TransDF. The results are presented in Table 6 and Table 6. Additional trends across different datasets and forecast lengths are provided in Appendix E.4. The primary observations are summarized as follows:
  • The coefficient $\alpha$ determines the relative importance of the transformed objective in (5). When $\alpha$ is set to 1, TransDF exclusively uses the transformed objective. We observe that increasing $\alpha$ from 0 to 1 generally leads to improved forecasting accuracy, with the best results typically achieved when $\alpha$ is close to 1. The performance improvement is significant, e.g., MSE reduction on ETTh2 by 0.016, showcasing the utility of the transformed objective to improve forecast performance.
  • The coefficient $\gamma$ determines the ratio of involved components for alignment. When $\gamma$ is set to 1, TransDF aligns all obtained components for model training. The results demonstrate that setting $\gamma$ to 1, with all label information preserved, does not necessarily yield optimal performance. Instead, the best results are often obtained at $\gamma<1$, rendering some loss of label information. For instance, $\gamma=0.7$ yields the best results on ETTm1 and ETTh2, while $\gamma=0.3$ is optimal for the Weather dataset. The rationale is that focusing on aligning the most significant components can reduce the task amount, thereby simplifying optimization. Since the majority of the information is contained in the most significant components, the information loss is minimal. Collectively, these factors contribute to improved forecast performance.
5 Conclusion
In this study, we highlight the importance of designing effective objectives for time-series forecasting. Two critical challenges are formulated: label autocorrelation, which induces bias, and the excessive number of tasks, which impedes optimization. To address these challenges, we introduce a model-agnostic learning objective called TransDF. This method transforms the label sequence into decorrelated components with discernible significance. Forecast models are trained to align the most significant components, which effectively mitigates label correlation due to the decorrelation between components and reduces task amount by discarding non-significant components. Experiments demonstrate that TransDF improves the performance of forecast models across diverse datasets.
Limitations & future works. In this work, we investigate the challenges of label correlation and exssive number of tasks in time-series forecasting. Nevertheless, these issues also manifest in areas such as speech generation, target recognition, and dense image prediction. Applying TransDF in these contexts is a promising avenue for future research. Additionally, historical sequence also exhibits autocorrelation and contains redundancy. Transforming inputs to derive decorrelated, compact representations could offer additional performance gains and also warrants investigation.
Appendix A On the implementation details of label correlation estimation
In this section, we introduce the motivation and implementation details of the label correlation estimation techniques in LABEL:fig:case. Measuring label correlation $Y_{t}\rightarrow Y_{t^{\prime}}$ is indeed challenging due to the presence of spurious correlations induced by the fork structure $Y_{t}\leftarrow L\rightarrow Y_{t^{\prime}}$ [28, 27]. This fork structure invalidates the use of Pearson correlation for accurately quantifying label correlation. To handle this problem, we follow our previous study [32], which involved the double machine learning (DML) method to estimate the ground-truth correlation while mitigating the fork variable’s influence.
The DML method is a statistical technique designed to estimate the causal effect of a treatment on an outcome while controlling for fork variables. Specifically, suppose we have a treatment variable $\mathcal{T}$, an outcome variable $\mathcal{Y}$, and a set of fork variables $\mathcal{X}$. The goal is to estimate the causal effect of $\mathcal{T}$ on $\mathcal{Y}$ while controlling for the influence of $\mathcal{X}$. To this end, DML first orthogonalizes both the treatment and outcome with respect to the fork variables. Two parametric models are employed to predict the treatment and outcome based on the fork variables. These predictions capture the impact of $\mathcal{X}$ on $\mathcal{Y}$ and $\mathcal{T}$. Subsequently, such impact of $\mathcal{X}$ is eliminated by calculating the residuals. Finally, the DML method regresses the outcome residuals on the treatment residuals, thereby measuring the causal effect of $\mathrm{T}$ on $\mathrm{Y}$ while removing the influence of the fork variables.
In our experiments, we measure label correlation by treating the input sequence $L$ as the fork variable and different steps of the label sequences $Y_{t}$ and $Y_{t^{\prime}}$ as the treatment and outcome variables, respectively. Then, we estimate the treatment effect of $Y_{t}$ on $Y_{t^{\prime}}$ controlling $L$. Similarly, when measuring the correlation between different components, we use different components $Z_{k}$ and $Z_{k^{\prime}}$ as the treatment and outcome variables. Linear regression model is employed as the parametric model for both the treatment and outcome variables for efficiency, which is consistent to [32].
To further complement the case study in LABEL:fig:auto, we analyzed the correlation matrices of the label sequences and the extracted components across multiple datasets, with the results presented in LABEL:fig:autoapp. The main observations are summarized as follows.
  • Panel (a) displays the correlation matrix of the label sequence, characterized by substantial non-diagonal elements, which highlight the strong autocorrelation among the labels. In contrast, panel (b) shows the correlation matrix of the extracted components, where the non-diagonal elements are nearly zero, indicating effective decorrelation.
  • Compared to the results reported in [32], where some obtained components remain correlated, the non-diagonal elements in panel (b) are fully eliminated. This difference arises because the Fourier transform in [32] achieves decorrelation only when the original label sequence is nearly infinitely long ($\mathrm{T}\rightarrow\infty$), a condition that is not met in real-world applications with finite forecast horizons. This limitation stems from the predefined nature of the projection matrix, which lacks adaptation to the specific properties of the data. In contrast, our method ensures decorrelation by solving a constrained optimization problem, without relying on an infinitely long forecast horizon, thereby providing a more reliable approach for handling autocorrelation bias.
Appendix B Theoretical Justification
Theorem B.1 (Autocorrelation bias, Theorem 3.1 in the main text).
Given label sequence $Y$ where $\Sigma\in\mathbb{R}^{\mathrm{T}\times\mathrm{T}}$ denotes the step-wise correlation coefficient, the TMSE in (1) is biased compared to the negative log-likelihood of the label sequence, which is given by:
$\mathrm{Bias}=\left\|Y-\hat{Y}\right\|_{\Sigma^{-1}}^{2}-\left\|Y-\hat{Y}\right\|^{2}-\frac{1}{2}\log\left|\Sigma\right|.$ (6)
where $\|Y-\hat{Y}\|_{\Sigma^{-1}}^{2}=(Y-\hat{Y})^{\top}\Sigma^{-1}(Y-\hat{Y})$. The bias vanishes if different steps in $Y$ are decorrelated. [^3] The pioneering work [32] identifies the bias under the first-order Markov assumption on the label sequence. This study generalizes this bias without the first-order Markov assumption.
Proof.
The proof is based on the properties of the multivariate normal distribution. It follows the flow in [32] but relaxes the first-order Markov assumption.
Suppose the label sequence follows a multivariate normal distribution with mean vector $\mu=[\hat{Y}_{1},\hat{Y}_{2},\ldots,\hat{Y}_{\mathrm{T}}]$ and covariance matrix $\Sigma$, where the off-diagonal entries are $\Sigma_{ij}=\rho_{ij}\sigma^{2}$ for $i\neq j$. Here, $\rho_{ij}$ denotes the partial correlation between $Y_{i}$ and $Y_{j}$ given the input sequence $L$. The log-likelihood of the label sequence $Y$ can be expressed as:
$\log p(Y)=\frac{1}{2}\left(\mathrm{T}\log(2\pi)+\log|\Sigma|+(Y-\hat{Y})^{\top}\Sigma^{-1}(Y-\hat{Y})\right).$
Removing the constant terms unrelated to $\hat{Y}$, we obtain the practical negative log-likelihood (PNLL):
$\mathrm{PNLL}=(Y-\hat{Y})^{\top}\Sigma^{-1}(Y-\hat{Y}).$
On the other hand, the TMSE loss can be expressed as:
$\mathrm{TMSE}=\left\|Y-\hat{Y}\right\|_{2}^{2}=(Y-\hat{Y})^{\top}I^{-1}(Y-\hat{Y}).$
where $I$ is the identity matrix. The difference between TMSE and PNLL can be expressed as:
$\mathrm{Bias}=\mathrm{PNLL}-\mathrm{TMSE}=(Y-\hat{Y})^{\top}\Sigma^{-1}(Y-\hat{Y})-(Y-\hat{Y})^{\top}I(Y-\hat{Y}),$
which immediately vanishes if the label sequence is decorrelated, i.e., $\Sigma=I$. The proof is completed. ∎
Lemma B.2 (Lemma 3.3 in the main text).
The projection matrix $\mathbf{P}^{*}$ can be obtained via singular value decomposition (SVD): $\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}(\mathbf{P}^{*})^{\top}$, where $\mathbf{U}\in\mathbb{R^{\mathrm{m\times m}}}$ and $\mathbf{P}^{*}\in\mathbb{R^{\mathrm{T\times T}}}$ consist of singular vectors, and the diagonal of $\mathbf{\Lambda}\in\mathbb{R^{\mathrm{m\times\mathrm{T}}}}$ consists of singular values.
Proof.
We first consider the case with $p=1$, where the orthogonal constrains are not involved. The Lagrangian can be written as:
$\mathcal{L}(\mathbf{P}_{1},\lambda_{1})=\mathbf{P}_{1}^{\top}\mathbf{S}\mathbf{P}_{1}-\lambda_{1}(\mathbf{P}_{1}^{\top}\mathbf{P}_{1}-1),$ (7)
where $\lambda_{1}$ is the Lagrangian multiplier. According to the first-order condition, the derivative with respect to $\mathbf{P}_{1}$ should be zero:
$\frac{\partial\mathcal{L}}{\partial\mathbf{P}_{1}}\Big{|}_{\mathbf{P}_{1}=\mathbf{P}_{1}^{*}}=2\mathbf{S}\mathbf{P}_{1}^{*}-2\lambda_{1}\mathbf{P}_{1}^{*}=0,$ (8)
which immediately follows by $\mathbf{S}\mathbf{P}_{1}^{*}=\lambda_{1}\mathbf{P}_{1}^{*}$. Apparently, $\mathbf{P}_{1}^{*}$ is an eigenvector of $\mathbf{S}$, with corresponding eigenvalue $\lambda_{1}$. Moreover, $\mathbf{P}_{1}^{*}$ is the leading eigenvector of $\mathbf{S}$ associated with the largest eigenvalue, which follows from the maximization objective is $\mathbf{P}_{1}{{}^{*}\top}\mathbf{S}\mathbf{P}_{1}^{*}=\lambda_{1}$,
We further consider the case with $p\geq 2$, impose orthogonality to all previous projection vectors. Defining Lagrangian multipliers $\lambda_{p}$ and $\{\mu_{j}\}_{j=1}^{p-1}$, we write the Lagrangian as follow
$\mathcal{L}(\mathbf{P}_{p},\lambda_{p},\{\mu_{j}\})=\mathbf{P}_{p}^{\top}\mathbf{S}\mathbf{P}_{p}-\lambda_{p}(\mathbf{P}_{p}^{\top}\mathbf{P}_{p}-1)-\sum_{j=1}^{p-1}\mu_{j}\mathbf{P}_{p}^{\top}\mathbf{P}_{j}.$ (9)
According to the first-order condition, the derivative with respect to $\mathbf{P}_{p}$ should be zero:
$\frac{\partial\mathcal{L}}{\partial\mathbf{P}_{p}}\Big{|}_{\mathbf{P}_{p}=\mathbf{P}_{p}^{*}}=2\mathbf{S}\mathbf{P}_{p}^{*}-2\lambda_{p}\mathbf{P}_{p}^{*}-\sum_{j=1}^{p-1}\mu_{j}\mathbf{P}_{j}=0.$ (10)
To resolve $\{\mu_{j}\}$, we take the inner product of both sides using $\mathbf{P}_{k}^{*\top}$ with $k=1,2,...,p-1$:
$\mathbf{P}_{k}^{*\top}\mathbf{S}\mathbf{P}_{p}^{*}-\lambda_{p}\mathbf{P}_{k}^{*\top}\mathbf{P}_{p}^{*}-\frac{1}{2}\sum_{j<p}\mu_{j}\mathbf{P}_{k}^{*\top}\mathbf{P}_{j}=0.$ (11)
Since $\mathbf{P}_{k}^{*}$ is previously obtained that satisfies $\mathbf{S}\mathbf{P}_{k}^{*}=\lambda_{k}\mathbf{P}_{k}^{*}$ and $\mathbf{P}_{k}^{*\top}\mathbf{P}_{k}^{*}=1$, we have $\mathbf{P}_{k}^{*\top}\mathbf{S}\mathbf{P}_{p}^{*}=\lambda_{k}$. Due to the orthogonal constraint, the current projection vector should be orthogonal to the previously derived ones, we have $\mathbf{P}_{k}^{*\top}\mathbf{P}_{p}^{*}=0$ and $\mathbf{P}_{k}^{*\top}\mathbf{P}_{j}=\delta_{k,j}$, where $\delta_{k,j}=1$ if $j=k$ and 0 otherwise. Putting together, the equation simplifies to: $\mu_{k}=0$.
Since $\mu_{k}=0$ holds for all $k=1,2,...,p-1$, we have $\mu_{1}=\mu_{2}=...=\mu_{p-1}=0$. Plugging back to (10), the optimal condition becomes:
$\mathbf{S}\mathbf{P}_{p}^{*}=\lambda_{p}\mathbf{P}_{p}^{*},$ (12)
with the additional restriction that $\mathbf{P}_{p}^{*}$ is orthogonal to all previous directions. Therefore, $\mathbf{P}_{p}^{*}$ must be the eigenvector of $\mathbf{S}$ corresponding to the $p$-th largest eigenvalue. That is, $\mathbf{P}$ can be derived by performing eigenvector decomposition on $\mathbf{S}$.
Moving forward, consider SVD: $\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top}$, we can represent $\mathbf{S}$ as
$\mathbf{S}=\mathbf{Y}^{\top}\mathbf{Y}=\mathbf{V}\mathbf{\Lambda}^{\top}\mathbf{U}^{\top}\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top}=\mathbf{V}\mathbf{\Lambda}^{2}\mathbf{V}^{\top},$ (13)
which implies that each column of $\mathbf{V}$ is the eigenvector of $\mathbf{S}$, i.e., $\mathbf{V}=\mathbf{P}^{*}$. Therefore, the projection matrix $\mathbf{P}^{*}$ can be obtained by performing SVD on $\mathbf{Y}$ as: $\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top}$, where $\mathbf{V}$ is exactly the optimum projection matrix $\mathbf{P}^{*}$. The proof is therefore completed. ∎
Lemma B.3 (Decorrelated components, Lemma 3.2 in the main text).
Suppose $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$ contains normalized label sequences for $\mathrm{m}$ samples, $\mathbf{Z}=[\mathbf{Z}_{1},...,\mathbf{Z}_{\mathrm{T}}]$ are the obtained components; for any $p\neq p^{\prime}$, we have $\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=0$.
Proof.
For any two latent components $\mathbf{Z}_{p}$ and $\mathbf{Z}_{p^{\prime}}$ with $p\neq p^{\prime}$, we have:
$\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=(\mathbf{Y}\mathbf{P}_{p})^{\top}(\mathbf{Y}\mathbf{P}_{p^{\prime}})=\mathbf{P}_{p}^{\top}\mathbf{Y}^{\top}\mathbf{Y}\mathbf{P}_{p^{\prime}}$ (14)
According to (12), $\mathbf{P}_{p}$ and $\mathbf{P}_{p^{\prime}}$ are eigenvectors of $\mathbf{Y}^{\top}\mathbf{Y}$, we have
$\mathbf{Y}^{\top}\mathbf{Y}\mathbf{P}_{p}=\lambda_{p}\mathbf{P}_{p},\qquad\mathbf{Y}^{\top}\mathbf{Y}\mathbf{P}_{p^{\prime}}=\lambda_{p^{\prime}}\mathbf{P}_{p^{\prime}},$ (15)
which immediately follows by $\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=\lambda_{p^{\prime}}\mathbf{P}_{p}^{\top}\mathbf{P}_{p^{\prime}}$. Recalling that different projection bases are constrained to orthogonal, i.e., $\mathbf{P}_{p}^{\top}\mathbf{P}_{p^{\prime}}=0$ for $p\neq p^{\prime}$, we have
$\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=0\qquad\text{for all }p\neq p^{\prime}.$ (16)
The proof is completed. ∎
Appendix C Generalized Orthogonalization Methods
In this section, we introduce alternative transform methods for obtaining latent components, each with distinct characteristics such as dimensionality reduction and noise isolation. We discuss their implications for transforming label sequences in time-series forecasting, with a comparative study detailed in Section 4.5.
RPCA.
The robust principal component analysis decomposes the data into a low-rank informative component and a sparse noise component, effectively separating structured signals from noise. Specifically, given $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$, it is achieved by solving:
$\min_{\mathbf{V},\mathbf{S}}\|\mathbf{V}\|_{*}+\lambda\|\mathbf{S}\|_{1},\quad\text{subject to}\quad\mathbf{Y}=\mathbf{V}+\mathbf{S},$ (17)
where $\|\cdot\|_{*}$ is the nuclear norm, $\|\cdot\|_{1}$ is the element-wise $\ell_{1}$ norm, and $\lambda$ is a regularization parameter. Afterwards, it performs the principal component analysis on the obtained informative component $\mathbf{V}$ to derive the projection matrix $\mathbf{P}$. The latent components are generated by $\mathbf{Z}=\mathbf{Y}\mathbf{P}$. While this approach enhances noise elimination, it does not guarantee decorrelation of the derived components, as the projection matrix $\mathbf{P}$ is derived from $\mathbf{V}$ instead of the original data matrix $\mathbf{Y}$.
SVD.
The singular value decomposition provides a method to decompose the matrix into different components. Given $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$, we have:
$\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top},$ (18)
where $\mathbf{U}\in\mathbb{R}^{\mathrm{m}\times\mathrm{r}}$ and $\mathbf{V}\in\mathbb{R}^{\mathrm{T}\times\mathrm{R}}$ are singular vectors, $\mathbf{\Lambda}\in\mathbb{R}^{\mathrm{r\times r}}$ is diagonal with rank $\mathrm{r}$. The right singular vector is used as the projection matrix, and the latent components are generated by $\mathbf{Z}=\mathbf{Y}\mathbf{V}$. One key distinction here needs to be highlighted. Unlike the workflow in the main text (Algorithm 1), the label sequence is not normalized after window generation before computing SVD here, resulting in non-decorrelated components.
FA.
Factor analysis models the observed data as linear combinations of a small number of latent factors plus noise, capturing the covariance structure through these unobserved factors. Specifically, given mean-centered $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$, the model assumes:
$\mathbf{Y}=\mathbf{V}\mathbf{F}^{\top}+\mathbf{E},$ (19)
where $\mathbf{V}\in\mathbb{R}^{\mathrm{m}\times\mathrm{K}}$ is the factor loading matrix, $\mathrm{K}$ is the number of latent factors ($\mathrm{K}\ll\mathrm{m}$), $\mathbf{F}\in\mathbb{R}^{\mathrm{T}\times\mathrm{K}}$ contains the latent factor scores for each sample, and $\mathbf{E}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$ is the noise matrix. The standard assumption is that each factor $f_{i}\sim\mathcal{N}(0,\mathbf{I})$ and noise $\epsilon_{i}\sim\mathcal{N}(0,\Psi)$, where $\Psi$ is a diagonal covariance matrix. The loadings $\mathbf{V}$ and factor scores $\mathbf{F}$ are typically estimated via maximum likelihood. The latent components are given by the estimated factor scores, i.e., $\mathbf{Z}=\mathbf{Y}\Psi^{-1}\mathbf{F}(\mathbf{I}+\mathbf{F}^{\top}\Psi^{-1}\mathbf{F})^{-1}:=\mathbf{Y}\mathbf{P}$ [^4] Adapted from source code of sklearn: https://github.com/scikit-learn/scikit-learn/blob/98ed9dc73/sklearn/decomposition/. This approach captures the covariance structure of $\mathbf{Y}$ via a small number of factors, but does not necessarily guarantee uncorrelated or noise-isolated components.
Appendix D Reproduction Details
D.1 Dataset descriptions
Table 7: Dataset description.
Dataset D Forecast length Train / validation / test Frequency Domain
ETTh1 7 96, 192, 336, 720 8545/2881/2881 Hourly Health
ETTh2 7 96, 192, 336, 720 8545/2881/2881 Hourly Health
ETTm1 7 96, 192, 336, 720 34465/11521/11521 15min Health
ETTm2 7 96, 192, 336, 720 34465/11521/11521 15min Health
Weather 21 96, 192, 336, 720 36792/5271/10540 10min Weather
ECL 321 96, 192, 336, 720 18317/2633/5261 Hourly Electricity
Traffic 862 96, 192, 336, 720 12185/1757/3509 Hourly Transportation
PEMS03 358 12, 24, 36, 48 15617/5135/5135 5min Transportation
PEMS08 170 12, 24, 36, 48 10690/3548/265 5min Transportation
  • Note: D denotes the number of variates. Frequency denotes the sampling interval of time points. Train, Validation, Test denotes the number of samples employed in each split. The taxonomy aligns with [35].
The datasets utilized in this study cover a wide range of time series data, detailed in Table 7, each exhibiting unique characteristics and temporal resolutions:
  • ETT [15] comprises data on 7 factors related to electricity transformers, collected from July 2016 to July 2018. This dataset is divided into four subsets: ETTh1 and ETTh2, with hourly recordings, and ETTm1 and ETTm2, documented every 15 minutes.
  • Weather [36] includes 21 meteorological variables gathered every 10 minutes throughout 2020 from the Weather Station of the Max Planck Biogeochemistry Institute.
  • ECL (Electricity Consumption Load) [36] presents hourly electricity consumption data for 321 clients.
  • Traffic [36] features hourly road occupancy rates from 862 sensors in the San Francisco Bay area freeways, spanning from January 2015 to December 2016.
  • PEMS [18] contains the public traffic network data in California collected by 5-minute windows. Two public subsets (PEMS03, PEMS08) are adopted in this work.
The datasets are chronologically divided into training, validation, and test sets, following the protocols outlined in [24, 19]. The dropping-last trick is disabled during the test phase. The length of the input sequence is standardized at 96 across the ETT, Weather, ECL, and Traffic datasets, with label sequence lengths set to 96, 192, 336, and 720. For the PEMS datasets, we employ label sequence lengths of 12, 24, 36, and 48.
D.2 Implementation Details
The baseline models in this study are reproduced using training scripts from the iTransformer repository [19] and Fredformer [23], after verifying reproducibility. Models are trained using the Adam optimizer [10], with learning rates chosen from the set $\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5}\}$ to minimize the MSE loss. Training is limited to a maximum of 10 epochs, with an early stopping mechanism triggered if there is no improvement in validation performance for 3 consecutive epochs.
In experiments where TransDF is integrated to enhance an existing forecast model, we adhere to the hyperparameter settings specified in the public benchmarks [19, 23], tuning only $\alpha$ within $[0,1]$, $\gamma$ within $(0,1]$, and the learning rate conservatively. Fine-tuning the learning rate is essential to account for the different magnitudes of the temporal and transformed losses. This fine-tuning is performed to minimize the MSE for different forecast lengths on the validation dataset.
Appendix E More Experimental Results
E.1 Complexity analysis
In this section, we analyze the running cost of TransDF. The core computation of TransDF involves (a) calculating the projection matrix $\mathbf{P}^{*}$ via SVD, and (b) performing transformation on both predicted and label sequences, followed by calculating their point-wise MAE loss. Given the target matrix $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$, the SVD step decomposes $\mathbf{Y}$ with an established complexity of $\mathcal{O}(\mathrm{m}\mathrm{T}^{2})$ (assuming $\mathrm{m}\geq\mathrm{T}$). For the sequence transformation, each sample (row) in $\mathbf{Y}$ is multiplied by the projection matrix $\mathbf{P}^{*}\in\mathbb{R}^{\mathrm{T}\times\mathrm{T}}$, resulting in a total complexity of $\mathcal{O}(\mathrm{m}\mathrm{T}^{2})$. The computation of point-wise MAE loss across all samples and forecast steps is $\mathcal{O}(\mathrm{m}\mathrm{T})$, which is negligible compared to the complexity of previous steps. Thus, the overall complexity per batch is dominated by the SVD and projection operations, both scaling as $\mathcal{O}(\mathrm{m}\mathrm{T}^{2})$. The main findings from the empirical evaluations are as follows.
  • LABEL:fig:running_cost (a) presents the computational cost for calculating the projection matrix. Overall, it increases linearly with the sample size and quadratically with the prediction length, which aligns with the theoretical complexity. Importantly, this operation is performed only once before training begins, rendering the associated overhead acceptable.
  • LABEL:fig:running_cost (b) presents the computational cost for the sequence transformation. The cost increases quadratically with the prediction length , but remains below 2 ms. This cost is comparable to that of a linear projection. Furthermore, sequence transformation is not required during inference.
In conclusion, TransDF does not add complexity to model inference, and the additional complexity during the training stage is negligible.
E.2 Overall performance
Table 8: The comprehensive results on the long-term forecasting task.
Models
PDF
Fredformer
iTransformer
FreTS
TimesNet
MICN
TiDE
DLinear
FEDformer
Autoformer
Transformer
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
96
0.321
0.357
0.326
0.361
0.338
0.372
0.342
0.375
0.368
0.394
0.319
0.366
0.353
0.374
0.346
0.373
0.401
0.434
0.485
0.468
0.503
0.482
192
0.360
0.378
0.365
0.382
0.382
0.396
0.385
0.400
0.406
0.409
0.364
0.395
0.391
0.393
0.380
0.390
0.415
0.446
0.504
0.482
0.807
0.664
336
0.389
0.400
0.396
0.404
0.427
0.424
0.416
0.421
0.454
0.444
0.395
0.425
0.423
0.414
0.413
0.414
0.432
0.450
0.520
0.489
0.847
0.678
720
0.447
0.435
0.459 0.444
0.496
0.463
0.513
0.489
0.527
0.474
0.505
0.499
0.486
0.448
0.472
0.450
0.522
0.500
0.594
0.523
1.037
0.771
Avg
0.379
0.393
0.387 0.398
0.411
0.414
0.414
0.421
0.438
0.430
0.396
0.421
0.413
0.407
0.403
0.407
0.442
0.457
0.526
0.491
0.799
0.648
ETTm2
96
0.172
0.251
0.177 0.260
0.182
0.265
0.188
0.279
0.184
0.262
0.178
0.277
0.182
0.265
0.188
0.283
0.205
0.289
0.218
0.300
0.386
0.441
192
0.235
0.294
0.242 0.300
0.257
0.315
0.264
0.329
0.257
0.308
0.266
0.343
0.247
0.304
0.280
0.356
0.271
0.334
0.282
0.340
1.410
0.881
336
0.293
0.333
0.302
0.340
0.320
0.354
0.322
0.369
0.315
0.345
0.299
0.354
0.307
0.343
0.375
0.420
0.327
0.366
0.335
0.370
1.940
1.070
720
0.388
0.389
0.399 0.397
0.423
0.411
0.489
0.482
0.452
0.421
0.489
0.482
0.408
0.398
0.526
0.508
0.428
0.425
0.423
0.420
2.914
1.276
Avg
0.272
0.317
0.280 0.324
0.295
0.336
0.316
0.365
0.302
0.334
0.308
0.364
0.286
0.328
0.342
0.392
0.308
0.354
0.315
0.358
1.662
0.917
ETTh1
96
0.368
0.391
0.377
0.396
0.385
0.405
0.398
0.409
0.399
0.418
0.381
0.416
0.387
0.395
0.389
0.404
0.391
0.433
0.449
0.465
1.028
0.778
192
0.424
0.422
0.437
0.425
0.440
0.437
0.451
0.442
0.452
0.451
0.497
0.489
0.439
0.425
0.442
0.440
0.418
0.448
0.459
0.469
1.010
0.789
336
0.467
0.441
0.486
0.449
0.480
0.457
0.501
0.472
0.488
0.469
0.589
0.555
0.482
0.447
0.488
0.467
0.487
0.484
0.511
0.500
0.908
0.743
720
0.465
0.463
0.488
0.467
0.504
0.492
0.608
0.571
0.549
0.515
0.665
0.617
0.484
0.471
0.505
0.502
0.494
0.514
0.488
0.498
0.987
0.785
Avg
0.431
0.429
0.447 0.434
0.452
0.448
0.489
0.474
0.472
0.463
0.533
0.519
0.448
0.435
0.456
0.453
0.447
0.470
0.477
0.483
0.983
0.774
ETTh2
96
0.282
0.330
0.293
0.344
0.301
0.349
0.315
0.374
0.321
0.358
0.351
0.398
0.291 0.340
0.330
0.383
0.351
0.391
0.355
0.397
1.485
0.959
192
0.359
0.381
0.372 0.391
0.383
0.397
0.466
0.467
0.418
0.417
0.492
0.489
0.376
0.392
0.439
0.450
0.456
0.456
0.478
0.471
4.218
1.585
336
0.394
0.414
0.420
0.433
0.425
0.432
0.522
0.502
0.464
0.454
0.656
0.582
0.417 0.427
0.589
0.538
0.477
0.492
0.459
0.469
2.775
1.361
720
0.400
0.427
0.421 0.439
0.436
0.448
0.792
0.643
0.434
0.450
0.981
0.718
0.429
0.446
0.757
0.626
0.522
0.505
0.499
0.502
2.274
1.257
Avg
0.359
0.388
0.377
0.402
0.386
0.407
0.524
0.496
0.409
0.420
0.620
0.546
0.378
0.401
0.529
0.499
0.452
0.461
0.448
0.460
2.688
1.291
ECL
96
0.145
0.235
0.161
0.258
0.150 0.242
0.180
0.266
0.170
0.272
0.170
0.281
0.197
0.274
0.197
0.282
0.187
0.302
0.189
0.304
0.253
0.350
192
0.159
0.249
0.174
0.269
0.168 0.259
0.184
0.272
0.183
0.282
0.185
0.297
0.197
0.277
0.197
0.286
0.207
0.322
0.271
0.371
0.262
0.356
336
0.173
0.264
0.194
0.290
0.182 0.274
0.199
0.290
0.203
0.302
0.190
0.298
0.212
0.292
0.209
0.301
0.211
0.326
0.243
0.352
0.269
0.363
720
0.203
0.292
0.235
0.319
0.214 0.304
0.234
0.322
0.294
0.366
0.221
0.329
0.254
0.325
0.245
0.334
0.253
0.361
0.295
0.388
0.277
0.365
Avg
0.170
0.260
0.191
0.284
0.179 0.270
0.199
0.288
0.212
0.306
0.192
0.302
0.215
0.292
0.212
0.301
0.214
0.328
0.249
0.354
0.265
0.358
Traffic
96
0.393
0.265
0.461
0.327
0.397 0.271
0.531
0.323
0.590
0.316
0.498
0.298
0.646
0.386
0.649
0.397
0.588
0.367
0.575
0.356
0.689
0.396
192
0.410
0.275
0.470
0.326
0.416 0.279
0.519
0.321
0.624
0.336
0.521
0.309
0.599
0.362
0.598
0.371
0.613
0.377
0.647
0.394
0.710
0.388
336
0.421
0.280
0.492
0.338
0.429 0.286
0.529
0.327
0.641
0.345
0.529
0.314
0.606
0.363
0.605
0.373
0.640
0.398
0.694
0.446
0.687
0.366
720
0.451
0.298
0.521
0.353
0.462 0.303
0.573
0.346
0.670
0.356
0.567
0.326
0.643
0.383
0.646
0.395
0.718
0.450
0.731
0.468
0.681
0.366
Avg
0.419
0.280
0.486
0.336
0.426 0.285
0.538
0.330
0.631
0.338
0.529
0.312
0.624
0.373
0.625
0.384
0.640
0.398
0.662
0.416
0.692
0.379
Weather
96
0.169
0.219
0.180
0.220
0.171
0.210
0.174
0.228
0.183
0.229
0.179
0.244
0.192
0.232
0.194
0.253
0.235
0.310
0.233
0.306
0.423
0.448
192
0.210
0.258
0.222
0.258
0.246
0.278
0.213
0.266
0.242
0.276
0.242
0.310
0.240
0.270
0.238
0.296
0.295
0.353
0.286
0.347
0.664
0.585
336
0.259
0.297
0.283
0.301
0.296
0.313
0.270
0.316
0.293
0.312
0.273
0.330
0.292
0.307
0.282
0.332
0.364
0.397
0.346
0.385
0.848
0.686
720
0.327
0.349
0.358
0.348
0.362
0.353
0.337
0.362
0.366
0.361
0.360
0.399
0.364
0.353
0.347
0.385
0.411
0.429
0.412
0.420
0.861
0.685
Avg
0.241
0.280
0.261
0.282
0.269
0.289
0.249
0.293
0.271
0.295
0.264
0.321
0.272
0.291
0.265
0.317
0.326
0.372
0.319
0.365
0.699
0.601
PEMS03
12
0.070
0.176
0.081
0.191
0.072 0.179
0.085
0.198
0.094
0.201
0.096
0.217
0.117
0.226
0.105
0.220
0.108
0.229
0.233
0.366
0.106
0.206
24
0.087
0.198
0.121
0.240
0.104
0.217
0.129
0.244
0.116
0.221
0.095 0.210
0.233
0.322
0.183
0.297
0.131
0.255
0.405
0.485
0.117
0.221
36
0.105
0.219
0.180
0.292
0.137
0.251
0.173
0.286
0.134
0.237
0.107 0.223
0.379
0.418
0.258
0.361
0.159
0.285
0.327
0.415
0.127
0.233
48
0.124
0.238
0.201
0.316
0.174
0.285
0.207
0.315
0.161
0.262
0.125 0.242
0.535
0.516
0.319
0.410
0.209
0.331
0.679
0.634
0.139
0.245
Avg
0.097
0.208
0.146
0.260
0.122
0.233
0.149
0.261
0.126
0.230
0.106 0.223
0.316
0.370
0.216
0.322
0.152
0.275
0.411
0.475
0.122
0.226
PEMS08
12
0.081
0.183
0.091
0.199
0.084 0.187
0.096
0.205
0.111
0.208
0.161
0.274
0.121
0.233
0.113
0.225
0.163
0.258
0.232
0.334
0.204
0.232
24
0.117
0.218
0.138
0.245
0.123 0.227
0.151
0.258
0.139
0.232
0.127
0.237
0.232
0.325
0.199
0.302
0.197
0.288
0.545
0.550
0.232
0.251
36
0.157
0.253
0.199
0.303
0.170
0.268
0.203
0.303
0.168 0.260
0.148
0.252
0.376
0.427
0.295
0.371
0.241
0.326
0.379
0.436
0.246
0.263
48
0.207 0.294
0.255
0.338
0.218
0.306
0.247
0.334
0.189
0.272
0.175
0.270
0.543
0.527
0.389
0.429
0.302
0.375
0.531
0.502
0.278
0.297
Avg
0.141
0.237
0.171
0.271
0.149
0.247
0.174
0.275
0.152
0.243
0.153
0.258
0.318
0.378
0.249
0.332
0.226
0.312
0.422
0.456
0.240
0.261
$1^{\text{st}}$ Count
43
42
0
1
0
1
0
0
1 1 1
0
0
0
0
0
0
0
0
0
0
0
  • Note: We fix the input length as 96 following [19]. Bold typeface highlights the top performance for each metric, while underlined text denotes the second-best results. Avg indicates the results averaged over forecasting lengths: T=96, 192, 336 and 720.
Table 9: The comprehensive results on the short-term forecasting task.
Models
PDF
Fredformer
iTransformer
FreTS
MICN
DLinear
Fedformer
Metric
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
Yearly
13.485
3.010
0.791
13.509 3.028 0.794
13.797
3.143
0.818
13.576
3.068
0.801
14.594
3.392
0.873
14.307
3.094
0.827
13.648
3.089
0.806
Quarterly
10.105
1.180
0.889
10.140 1.185 0.893
10.503
1.248
0.932
10.361
1.223
0.916
11.417
1.385
1.023
10.500
1.237
0.928
10.612
1.246
0.936
Monthly
12.649
0.930
0.875
12.696 0.931 0.878
13.227
1.013
0.935
13.088
0.990
0.919
13.834
1.080
0.987
13.362
1.007
0.937
14.181
1.105
1.011
Others
4.852
3.274
1.027
4.848
3.230
1.019
5.101
3.419
1.076
5.563
3.71
1.17
6.137
4.201
1.308
5.12
3.649
1.114
4.823
3.243 1.019
Average
11.841
1.585
0.851
11.879 1.590 0.854
12.298
1.68
0.893
12.169
1.66
0.883
13.044
1.841
0.962
12.48
1.674
0.898
12.734
1.702
0.914
$1^{\text{st}}$ Count
4
4
4
0
1 1
0
0
0
0
0
0
0
0
0
0
0
0
1
0
0
  • Note: Bold typeface highlights the top performance for each metric, while underlined text denotes the second-best results. Avg indicates the results averaged over forecasting lengths: yearly, quarterly, and monthly.
Long-term Forecast.
A comprehensive performance comparison for long-term forecasting across various datasets is presented in Table 8. The best-performing baseline models for each dataset were utilized to implement the TransDF paradigm. Specifically, Fredformer was employed for the ETT datasets, iTransformer for the ECL, Traffic, and PEMS08 datasets, FreTS for the Weather dataset, and MICN for the PEMS03 dataset. TransDF consistently enhanced the performance of each base model, achieving the lowest MSE in 43 out of 45 cases and the lowest MAE in 42 out of 45 cases. The rare exceptions where TransDF did not achieve the lowest MSE were attributed to the inherent superiority of other models over the base model in specific datasets and forecast lengths (e.g., iTransformer versus FreTS on the Weather dataset with $\mathrm{T}=96$).
Short-term Forecast.
The short-term forecasting task was evaluated in Table 9, with Fredformer [23] serving as the forecasting model in the TransDF implementation. Similar to the long-term forecasting results, TransDF improved Fredformer’s performance in most cases. However, there were three instances where Fredformer outperformed TransDF. This discrepancy occurred because the critical hyperparameters ($\alpha$ and $\gamma$) were tuned to minimize the average validation error across all forecast lengths, rather than being fine-tuned for each individual forecast length. Although fine-tuning $\alpha$ and $\gamma$ for each forecast length is possible, it was deemed unnecessary, as the current results sufficiently demonstrate TransDF’s effectiveness.
Showcases.
Additional examples of forecast sequences are provided in LABEL:fig:pred_app_ettm2_336 and LABEL:fig:pred_app_ecl_192. TransDF effectively captured the principal patterns and closely followed the label sequences. These successes are attributed to TransDF’s unique capability to operate in the transformed domain, where autocorrelation challenges are mitigated and the significance of important components is emphasized.
E.3 Generalization Studies
Varying forecast models.
In this section, we further investigate the versatility of TransDF in enhancing a variety of forecasting models, including Fredformer, iTransformer, FreTS, and DLinear. The results, presented in LABEL:fig:backbone_app, cover four distinct datasets and are averaged across forecast lengths (96, 192, 336, 720), with error bars representing 50% confidence intervals. TransDF consistently improves the performance of these forecasting models, with especially notable gains observed in Transformer-based architectures such as Fredformer and iTransformer. These findings confirm the effectiveness of TransDF in boosting neural forecasting models and underscore its potential as a versatile training methodology for time series forecasting.
Varying Transformations.
Table 10: Varying projection bases results.
Trans
PCA
RPCA
SVD
FA
DF
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ECL
96
0.1449
0.2348
0.1450 0.2349
0.1450
0.2350
0.1478
0.2385
0.1500
0.2415
192
0.1592
0.2487
0.1594 0.2487
0.1595
0.2490
0.1619
0.2517
0.1681
0.2591
336
0.1731 0.2645
0.1732
0.2646
0.1730
0.2643
0.1789
0.2711
0.1823
0.2744
720
0.2033
0.2920
0.2066 0.2960
0.2214
0.3066
0.2095
0.2975
0.2145
0.3035
Avg
0.1701
0.2600
0.1710 0.2611
0.1747
0.2637
0.1745
0.2647
0.1787
0.2696
Weather
96
0.1692
0.2185
0.1715 0.2199
0.1723
0.2223
0.1717
0.2247
0.1737
0.2277
192
0.2102
0.2575
0.2116
0.2590 0.2116
0.2597
0.2125
0.2636
0.2128
0.2661
336
0.2586
0.2971
0.2631 0.3072
0.2676
0.3110
0.2613
0.2997
0.2705
0.3159
720
0.3271
0.3487
0.3303 0.3581
0.3394
0.3681
0.3354
0.3610
0.3372
0.3623
Avg
0.2413
0.2805
0.2441 0.2860
0.2477
0.2903
0.2452
0.2872
0.2486
0.2930
In this section, we provide a detailed performance comparison across different forecast lengths and transformation methods on the ECL and Weather datasets. The results, shown in Table 10, indicate that TransDF consistently and significantly improves the performance of DF, regardless of the orthogonal transformation employed. Notably, label alignment with principal component awareness yields the best results across almost all datasets and forecast lengths. Although, on the ECL dataset with $\mathrm{T}=336$, the SVD method achieves the lowest MSE and MAE, the PCA-based TransDF delivers comparable performance with lighter computational requirements. These results highlight the importance of orthogonal label alignment within TransDF, which is crucial for eliminating autocorrelation.
E.4 Hyperparameter Sensitivity
In this section, we examine how adjusting the transformed loss strength $\alpha$ and the projection rank ratio $\gamma$ impacts the performance of TransDF across six datasets: ETTm1, ETTm2, ETTh1, ETTh2, ECL, and Weather. Fredformer is used as the forecasting model. The MSE results are presented in LABEL:fig:sensi-alpha and LABEL:fig:sensi-rank.
Varying transformed loss strength.
We find that increasing $\alpha$ from 0 to 1 generally reduces the forecast error across different datasets and forecast lengths, highlighting the benefits of learning in the orthogonal domain. Notably, the minimum forecast error often occurs at $\alpha$ values close to 1 rather than exactly at 1; for instance, the optimal value for the ETTh1 dataset is 0.8. This suggests that integrating supervisory signals from both the time and orthogonal domains can enhance forecasting performance. However, the improvement may be incremental compared to simply setting $\alpha=1$.
Varying ratio of projection rank.
We generally observe that increasing $\gamma$ from 0 to 1 first reduces the forecast error and then increases it, with the minimum error typically occurring at a ratio close to 1. For example, the minimum forecast error is achieved at $\gamma=0.8$ for both the ETTh1 and ECL datasets. Notably, in some cases—such as with the Weather dataset—a low rank ratio consistently outperforms full rank, suggesting that full-rank alignment may introduce redundancy. These results underscore the importance of principal component awareness within TransDF and suggest that low-rank projection can be more beneficial than full-rank alignment.
E.5 Comparison with Different Label Alignment Strategies
Table 11: Comparable results with different label alignment strategies and Fredformer.
Loss TransDF FreDF Koopman Dilate Soft-DTW DPTA DF
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.321 0.357 0.326 0.355 0.335 0.368 0.337 0.367 0.332 0.363 0.332 0.364 0.326 0.361
192 0.360 0.378 0.363 0.380 0.366 0.384 0.364 0.384 0.370 0.386 0.370 0.386 0.365 0.382
336 0.389 0.400 0.392 0.400 0.399 0.408 0.397 0.406 0.406 0.409 0.409 0.410 0.396 0.404
720 0.447 0.435 0.455 0.440 0.456 0.441 0.457 0.443 0.478 0.450 0.476 0.448 0.459 0.444
Avg 0.379 0.393 0.384 0.394 0.389 0.400 0.389 0.400 0.397 0.402 0.396 0.402 0.387 0.398
ETTh1 96 0.368 0.391 0.370 0.392 0.375 0.397 0.378 0.399 0.376 0.398 0.378 0.399 0.377 0.396
192 0.424 0.422 0.436 0.437 0.438 0.434 0.439 0.435 0.439 0.435 0.438 0.433 0.437 0.425
336 0.467 0.441 0.473 0.443 0.473 0.455 0.481 0.453 0.484 0.455 0.486 0.455 0.486 0.449
720 0.465 0.463 0.474 0.466 0.523 0.487 0.516 0.482 0.542 0.510 0.538 0.510 0.488 0.467
Avg 0.431 0.429 0.438 0.434 0.452 0.443 0.453 0.442 0.460 0.449 0.460 0.449 0.447 0.434
ECL 96 0.151 0.245 0.152 0.247 0.166 0.263 0.158 0.253 0.168 0.266 0.158 0.253 0.161 0.258
192 0.166 0.256 0.166 0.257 0.174 0.267 0.170 0.263 0.218 0.313 0.216 0.307 0.174 0.269
336 0.181 0.274 0.183 0.278 0.188 0.280 0.190 0.286 0.197 0.291 0.199 0.295 0.194 0.290
720 0.213 0.304 0.216 0.304 0.232 0.318 0.229 0.316 0.240 0.322 0.235 0.322 0.235 0.319
Avg 0.178 0.270 0.179 0.272 0.190 0.282 0.187 0.280 0.206 0.298 0.202 0.294 0.191 0.284
Weather 96 0.171 0.208 0.174 0.213 0.174 0.214 0.173 0.214 0.173 0.213 0.179 0.219 0.180 0.220
192 0.219 0.253 0.219 0.254 0.220 0.256 0.225 0.260 0.220 0.255 0.223 0.257 0.222 0.258
336 0.277 0.295 0.278 0.296 0.280 0.298 0.280 0.299 0.281 0.296 0.281 0.298 0.283 0.301
720 0.353 0.346 0.354 0.347 0.354 0.347 0.355 0.348 0.369 0.355 0.356 0.347 0.358 0.348
Avg 0.255 0.276 0.256 0.277 0.257 0.279 0.258 0.280 0.261 0.280 0.260 0.280 0.261 0.282
Table 12: Comparable results with different label alignment strategies and iTransformer.
Loss TransDF FreDF Koopman Dilate Soft-DTW DPTA DF
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.323 0.358 0.334 0.365 0.350 0.382 0.342 0.376 0.339 0.373 0.341 0.375 0.338 0.372
192 0.371 0.388 0.381 0.390 0.389 0.400 0.381 0.396 0.383 0.395 0.383 0.395 0.382 0.396
336 0.408 0.407 0.417 0.412 0.425 0.423 0.418 0.418 0.429 0.423 0.429 0.423 0.427 0.424
720 0.477 0.450 0.489 0.453 0.489 0.458 0.487 0.457 0.516 0.469 0.512 0.467 0.496 0.463
Avg 0.395 0.401 0.405 0.405 0.413 0.416 0.407 0.412 0.417 0.415 0.416 0.415 0.411 0.414
ETTh1 96 0.378 0.393 0.378 0.395 0.392 0.411 0.385 0.405 0.387 0.405 0.386 0.405 0.385 0.405
192 0.428 0.423 0.428 0.423 0.446 0.442 0.440 0.437 0.443 0.439 0.441 0.439 0.440 0.437
336 0.473 0.450 0.470 0.447 0.483 0.461 0.480 0.457 0.494 0.464 0.489 0.462 0.480 0.457
720 0.473 0.469 0.490 0.484 0.501 0.491 0.504 0.492 0.557 0.520 0.538 0.509 0.504 0.492
Avg 0.438 0.434 0.442 0.437 0.455 0.451 0.452 0.448 0.470 0.457 0.463 0.454 0.452 0.448
ECL 96 0.145 0.235 0.149 0.238 0.151 0.243 0.150 0.241 0.149 0.241 0.149 0.240 0.150 0.242
192 0.159 0.249 0.163 0.251 0.167 0.257 0.168 0.259 0.164 0.255 0.166 0.257 0.168 0.259
336 0.173 0.264 0.179 0.268 0.182 0.275 0.181 0.274 0.180 0.274 0.180 0.272 0.182 0.274
720 0.203 0.292 0.212 0.297 0.212 0.300 0.212 0.300 0.207 0.296 0.212 0.300 0.214 0.304
Avg 0.170 0.260 0.176 0.264 0.178 0.269 0.178 0.269 0.175 0.266 0.177 0.267 0.179 0.270
Weather 96 0.163 0.202 0.170 0.208 0.206 0.257 0.208 0.259 0.207 0.252 0.209 0.258 0.171 0.210
192 0.214 0.248 0.219 0.252 0.264 0.300 0.252 0.285 0.264 0.303 0.258 0.291 0.246 0.278
336 0.274 0.294 0.279 0.296 0.309 0.326 0.311 0.328 0.314 0.333 0.312 0.331 0.296 0.313
720 0.351 0.344 0.358 0.347 0.377 0.369 0.374 0.364 0.384 0.377 0.383 0.373 0.362 0.353
Avg 0.251 0.272 0.257 0.276 0.289 0.313 0.286 0.309 0.292 0.316 0.291 0.313 0.269 0.289
In this section, we present a comprehensive comparison between TransDF and methods that employ different learning objectives to align the forecast sequence with the label sequence: (1) FreDF [32] and Koopman [12], which use frequency domain alignment; and (2) Dilate [13], Soft-DTW [3], and DPTA [25], which utilize shape alignment approaches. For a fair comparison, we integrated the official implementations of these loss functions into both the Fredformer and iTransformer models. As shown in Table 11 and Table 12, TransDF consistently and significantly outperforms these methods across all datasets, model architectures, and forecast lengths. This superior performance can be attributed to TransDF’s unique ability to debias the learning objective in the orthogonal domain and focus on principal components—capabilities not offered by the other methods.
E.6 Case Study with PatchTST and Varying Input Lengths
Table 13: Varying input sequence length results on the Weather dataset.
Models
TransDF iTransformer TransDF PatchTST
Metrics
MSE MAE MSE MAE MSE MAE MSE MAE
Input sequence length 96 96 0.163 0.202 0.171 0.210 0.175 0.213 0.200 0.244
192 0.214 0.248 0.246 0.278 0.224 0.257 0.229 0.263
336 0.274 0.294 0.296 0.313 0.276 0.296 0.287 0.303
720 0.351 0.344 0.362 0.353 0.353 0.346 0.363 0.353
Avg 0.250 0.272 0.269 0.289 0.257 0.278 0.270 0.291
192 96 0.163 0.205 0.168 0.215 0.158 0.199 0.164 0.208
192 0.210 0.248 0.213 0.253 0.204 0.242 0.225 0.269
336 0.259 0.287 0.265 0.294 0.257 0.286 0.287 0.308
720 0.334 0.338 0.341 0.345 0.332 0.337 0.341 0.345
Avg 0.241 0.270 0.247 0.277 0.238 0.266 0.254 0.283
336 96 0.157 0.203 0.162 0.213 0.150 0.196 0.156 0.206
192 0.199 0.246 0.211 0.256 0.196 0.241 0.222 0.277
336 0.251 0.287 0.260 0.295 0.246 0.282 0.251 0.285
720 0.324 0.338 0.332 0.341 0.320 0.333 0.327 0.338
Avg 0.233 0.268 0.241 0.276 0.228 0.263 0.239 0.277
720 96 0.161 0.213 0.172 0.225 0.152 0.201 0.154 0.207
192 0.205 0.250 0.220 0.268 0.198 0.248 0.205 0.254
336 0.254 0.292 0.282 0.311 0.248 0.284 0.248 0.288
720 0.318 0.339 0.337 0.351 0.313 0.335 0.317 0.339
Avg 0.235 0.274 0.253 0.289 0.228 0.267 0.231 0.272
In this section, we focus on iTransformer [19] and PatchTST [21], highlighting the effectiveness of TransDF in enhancing their performance with varying input sequence lengths—complementing the fixed length of 96 used in the main text. As shown in Table 13, TransDF consistently improves the performance of both iTransformer and PatchTST across different input lengths. Notably, under our experimental settings, PatchTST with $\mathrm{H}=336$ achieves results comparable to the original “PatchTST/42” results reported in [21], while TransDF further reduces both the MSE and MAE by 0.01, demonstrating its robustness across different input lengths.
E.7 Random Seed Sensitivity
Table 14: Experimental results ($\mathrm{mean}_{\pm\mathrm{std}}$) with varying seeds (2021-2025).
Dataset
ECL Weather
Models TransDF DF TransDF DF
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
96 0.145±0.000 0.235±0.000 0.150±0.001 0.242±0.001 0.164±0.001 0.203±0.001 0.190±0.012 0.232±0.014
192 0.160±0.001 0.249±0.001 0.166±0.002 0.257±0.002 0.216±0.002 0.250±0.001 0.240±0.011 0.272±0.010
336 0.174±0.002 0.266±0.002 0.181±0.001 0.273±0.001 0.274±0.001 0.294±0.001 0.293±0.003 0.310±0.003
720 0.205±0.001 0.293±0.001 0.216±0.004 0.303±0.003 0.353±0.002 0.344±0.001 0.361±0.002 0.352±0.001
Avg 0.171±0.001 0.261±0.001 0.178±0.001 0.269±0.001 0.252±0.001 0.273±0.001 0.271±0.003 0.292±0.003
In this section, we investigate the sensitivity of the results to the choice of random seeds. To this end, we report the mean and standard deviation of the results obtained from experiments using five different random seeds (2021, 2022, 2023, 2024, 2025) in Table 14. We examine both (1) iTransformer and (2) TransDF, which is applied to refine iTransformer. The results on both the ECL and Weather datasets show minimal sensitivity to random seeds, with standard deviations below 0.005 in all averaged cases.