### v2

# Time-o1: Time-Series Forecasting Needs Transformed Label Alignment

Abstract
Training time-series forecast models presents unique challenges in designing effective learning objectives. Existing methods predominantly utilize the temporal mean squared error, which faces two critical challenges: (1) label autocorrelation, which leads to bias from the label sequence likelihood; (2) excessive amount of tasks, which increases with the forecast horizon and complicates optimization. To address these challenges, we propose Time-o1, a transformation-augmented learning objective tailored for time-series forecasting. The central idea is to transform the label sequence into decorrelated components with discriminated significance. Models are then trained to align the most significant components, thereby effectively mitigating label autocorrelation and reducing task amount. Extensive experiments demonstrate that Time-o1 achieves state-of-the-art performance and is compatible with various forecast models. Code is available at https://github.com/Master-PLC/Time-o1.
1 Introduction
Time-series forecasting involves predicting future data from historical observations qiu2025easytime ; k2vae and has been applied across diverse domains, such as weather forecasting in meteorology (application_weather, ), user behavior analysis in e-commerce (chenmode, ), and process monitoring in manufacturing (wang2025tnnlspot, ; wang2024taiattentionmixer, ). To build effective forecast models, there are two questions that warrant investigation: (1) How to design a neural network architecture to encode historical observations, and (2) How to devise a learning objective to train the neural network. Both are critical for model performance.
Recent research has primarily focused on developing neural network architectures. The key challenge lies in exploiting the autocorrelation in the historical sequences (wu2024autocts++, ). To this end, various architectures have been proposed (qiu2024tfb, ; li2024foundts, ; wu2024fully, ), such as recurrent neural networks (S4, ), convolutional neural networks (Timesnet, ; micn, ), and graph neural networks (FourierGNN, ). The current progress is marked by a debate between Transformers and simple linear models. Transformers, equipped with self-attention mechanisms, offer superior scalability (itransformer, ; PatchTST, ). In contrast, linear models, which encapsulate temporal dynamics using linear layers, are straightforward to implement and often demonstrate strong performance (DLinear, ). These advancements showcase the rapid evolution in neural architecture design for time-series forecasting.
In contrast, the design of learning objectives has received less attention wang2025iclrfredf ; dbloss ; psloss . Most existing methods employ the temporal mean squared error (TMSE) as the learning objective, which measures the step-wise discrepancy between forecast and label sequences (itransformer, ; PatchTST, ). While being effective for various scenarios, it has two critical limitations. First, it is biased against the true likelihood of label sequence due to the presence of autocorrelation in the label sequence. Second, the number of prediction tasks increases with the forecast horizon, which complicates the optimization process since multitask learning is known to be challenging given excessive tasks (mtlsurvey, ; cagrad, ). These challenges present unique challenges in designing learning objectives for time-series forecasting.
To handle these challenges, we propose a transformation-augmented learning objective tailored for time-series forecasting. The key idea is to transform the label sequence into decorrelated components ranked by significance. By aligning the most significant decorrelated components, Time-o1 mitigates label autocorrelation and reduces the number of tasks.
Our main contributions are summarized as follows:
  • We formulate two critical challenges in designing objectives for time-series forecasting: label autocorrelation that induces bias, and the excessive number of tasks that impedes optimization.
  • We propose Time-o1, which transforms labels into decorrelated components with discriminated significance. By aligning the significant components, it addresses the two challenges above.
  • We conduct comprehensive experiments to demonstrate Time-o1’s efficacy, consistently boosting the performance of state-of-the-art forecast models across diverse datasets.
2 Preliminaries
This paper focuses on the time-series forecasting problem qiu2025easytime . By the way of preface, uppercase letters (e.g., $Y$) denote random variables, and bolded letters (e.g., $\mathbf{Y}$) denote matrices containing data or parameters. One key distinction warrants emphasis: we are concentrating on the design of learning objectives for training forecast models wang2025iclrfredf ; le2020probabilistic ; le2019shape , rather than on the design of neural network architectures to implement the forecast models itransformer ; DLinear .
Suppose $X$ is a time-series dataset with $\mathrm{D}$ covariates, where $X_{n}$ denotes the observation at the $n$-th step. At an arbitrary $n$-th step, the historical sequence is defined as $L=[X_{n-\mathrm{H}+1},\ldots,X_{n}]\in\mathbb{R}^{\mathrm{H}\times\mathrm{D}}$, the label sequence is defined as $Y=[X_{n+1},\ldots,X_{n+\mathrm{T}}]\in\mathbb{R}^{\mathrm{T}\times\mathrm{D}}$, where $\mathrm{H}$ is the historical length and $\mathrm{T}$ is the forecast horizon. The target of time-series forecasting is to train a model $g:\mathbb{R}^{\mathrm{H}\times\mathrm{D}}\rightarrow\mathbb{R}^{\mathrm{T}\times\mathrm{D}}$ that generates accurate prediction sequence $\hat{Y}$ approximating the label sequence.
There are two aspects to building forecast models: (1) neural network architectures that effectively encode historical sequences, and (2) learning objectives for training these neural networks. While this paper focuses on the learning objective, we provide a brief review of both aspects for contextualization.
2.1 Model architectures for time-series forecasting
Neural networks have been pervasive in encoding historical sequences for their capability of automating feature interactions and capturing nonlinear correlations. Notable examples include RNNs (e.g., S4 (S4, ), Mamba), CNNs (e.g., TimesNet (Timesnet, )), and GNNs (e.g., MTGNN (Mtgnn, )), each tailored to encode the dynamics within input sequences. The current progress centers on the comparison between Transformer-based and MLP-based architectures. Transformers (e.g., PatchTST (PatchTST, ), iTransformer (itransformer, )) exhibit substantial scalability with increasing data size but entail high computational costs. In contrast, MLPs (e.g., DLinear (DLinear, ), TimeMixer (wang2024timemixer, )) are generally more efficient but less scalable with larger datasets and struggle to handle varying input lengths.
2.2 Learning objectives for time-series forecasting
Modern time-series models predominantly adopt the direct forecast paradigm, generating T-step forecasts simultaneously using a multi-output head Informer ; itransformer ; DLinear . The learning objective is typically the temporal mean squared error (TMSE) between the forecast and label sequences, given by:
$\mathcal{L}_{\mathrm{tmse}}=\sum_{t=1}^{\mathrm{T}}\left(Y_{t}-\hat{Y}_{t}\right)^{2},$ (1)
which is widely employed in recent studies (e.g., FreTS (FreTS, ), iTransformer (itransformer, ), FredFormer (fredformer, ), DUET qiu2025duet ). However, this objective has been shown to be biased due to the autocorrelation present in the label sequence wang2025iclrfredf . To address this bias, one line of research advocates for shape alignment between the forecast and label sequences to exploit autocorrelation (e.g., Dilate (le2019shape, ) and Soft-DTW (soft-dtw, )). However, these methods lack rigorous theoretical guarantees for unbiased objective and empirical evidence of improved performance. Another notable approach involves computing the forecast error in the frequency domain, which reduces bias with theoretical guarantees (wang2025iclrfredf, ).
3 Methodology
3.1 Motivation
The learning objective is a fundamental component in training effective forecast models, yet its importance remains underexplored. Existing approaches predominantly employ the TMSE in (1) as the objective (itransformer, ; PatchTST, ; OLinear, ). This practice, however, encounters two fundamental limitations rooted in the characteristics of time-series forecasting task.
First, TMSE introduces bias due to autocorrelation. In time-series forecasting, observations exhibit strong dependencies on their past values (DLinear, ), resulting in step-wise correlation in the label sequence. In contrast, TMSE treats the forecast of each step as an independent task, thereby neglecting these correlations. This mismatch makes TMSE biased with respect to the true likelihood of the label sequence, as presented in Theorem 3.1.
Theorem 3.1 (Autocorrelation bias).
Given label sequence $Y$ where $\Sigma\in\mathbb{R}^{\mathrm{T}\times\mathrm{T}}$ denotes the step-wise correlation coefficient, the TMSE in (1) is biased compared to the negative log-likelihood of the label sequence, which is given by:
$\mathrm{Bias}=\left\|Y-\hat{Y}\right\|_{\Sigma^{-1}}^{2}-\left\|Y-\hat{Y}\right\|^{2}-\frac{1}{2}\log\left|\Sigma\right|.$ (2)
where $\|v\|_{\Sigma^{-1}}^{2}=v^{\top}\Sigma^{-1}v$. The bias vanishes if different steps in $Y$ are decorrelated. [^1] The pioneering work (wang2025iclrfredf, ) identifies the bias under the first-order Markov assumption on the label sequence. This study generalizes this bias without the first-order Markov assumption.
Second, TMSE poses optimization difficulties as the forecast horizon grows. The large forecast horizon is crucial for applications such as manufacturing (enabling comprehensive production planning (wang2024taselsptd, ; wang2024tiispoti, )) and transportation (enabling proactive traffic management (application_traffic, )). As TMSE treats each forecasted step as an independent task, a large horizon results in excessive tasks. However, optimization is known to be difficult given excessive tasks (mtlsurvey, ; cagrad, ), as gradients from different tasks often conflict (mtlsurvey, ; cagrad, ), impeding convergence and leading to suboptimal model performance.
Designing effective learning objectives to handle the two limitations is challenging. The previous work FreDF (wang2025iclrfredf, ) proposes a frequency loss, which transforms the label and forecast sequences into frequency components and aligns them in the frequency domain. This approach is motivated by Theorem 3.1: bias vanishes if different components are decorrelated. However, the decorrelation of frequency components holds only when the forecast horizon $\mathrm{T}\rightarrow\infty$ (see Theorem 3.3 in (wang2025iclrfredf, )). In real-world settings with finite horizon, frequency components remain correlated, rendering FreDF ineffective in eliminating bias. Additionally, the optimization difficulty remains, since transforming to the frequency domain retains the label length. Consequently, FreDF does not fully address the autocorrelation bias and the optimization difficulty.
Given the critical role of objective in training forecast models and the limitations of existing methods, it is compelling to develop an innovative objective to address the limitations and advance forecast performance. Importantly, there are two questions that warrant investigation. How to devise an objective that eliminates autocorrelation and reduces task amount? Does it improve forecast performance?
(a) Comparison of label sequence and associated components. (a) shows the correlation volume within the label sequence (left panel) and components (right panel). (b) visualizes 8 label sequences randomly from ETTh1 (left panel) and the associated components (right panel).
3.2 Transforming label sequence with optimized projection matrix
In this section, we present a method for transforming label sequences into latent components to eliminate autocorrelation and distinguish significant components. Suppose $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$ contains normalized label sequences for $\mathrm{m}$ samples, $\mathbf{P}=\left[\mathbf{P}_{1},\mathbf{P}_{2},...,\mathbf{P}_{\mathrm{T}}\right]$ is the projection matrix, and the components are produced as $\mathbf{Z}=\mathbf{Y}\mathbf{P}$. The target is for $\mathbf{Z}$ to be decorrelated and ranked by significance. For example, FreDF specifies $\mathbf{P}$ as a Fourier matrix, which does not adapt to specific data properties and thus fails to decorrelate the components and distinguish the significant components [^2] In the subsequent paragraphs, we use the univariate case with $\mathrm{D}=1$ for clarity. In the multivariate case, different variates can be treated separately to produce decorrelated components..
A natural approach to obtaining the projection matrix $\mathbf{P}$ is solving optimization problem with constraints to ensure the desired properties. To find the $p$-th component, the projection vector can be calculated by solving the following problem:
$\mathbf{P}_{p}^{*}=$ $\ \text{arg}\underset{\mathbf{P}_{p}}{\text{max}}$ $\left(\mathbf{Y}\mathbf{P}_{p}\right)^{\top}\left(\mathbf{Y}\mathbf{P}_{p}\right)$ (3)
[…]
where $\mathbf{Z}_{p}=\mathbf{Y}\mathbf{P}_{p}$ is the $p$-th component, the normalization constraint $\|\mathbf{P}_{p}\|^{2}=1$ is imposed to avoid trivial solution: $\mathbf{P}_{p}\rightarrow\infty$. The optimization target is to maximize the variance of $\mathbf{Z}_{p}$, which is equivalent to maximizing its significance, as components with larger variance contain richer information. For $p>1$, the projection axis is required to be orthogonal to the previous axes to avoid redundancy. By solving the optimizations above from $p=1$ to $\mathrm{T}$ sequentially, we obtain the projection matrix $\mathbf{P}^{*}=\left[\mathbf{P}_{1}^{*},...,\mathbf{P}_{\mathrm{T}}^{*}\right]$. The components are then produced as $\mathbf{Z}=\mathbf{Y}\mathbf{P}^{*}$.
Lemma 3.2 (Decorrelated components).
Suppose $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$ contains normalized label sequences for $\mathrm{m}$ samples, $\mathbf{Z}=[\mathbf{Z}_{1},...,\mathbf{Z}_{\mathrm{T}}]$ are the obtained components; for any $p\neq p^{\prime}$, we have $\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=0$.
Lemma 3.3.
The projection matrix $\mathbf{P}^{*}$ can be obtained via singular value decomposition (SVD): $\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}(\mathbf{P}^{*})^{\top}$, where $\mathbf{U}\in\mathbb{R^{\mathrm{m\times m}}}$ and $\mathbf{P}^{*}\in\mathbb{R^{\mathrm{T\times T}}}$ consist of singular vectors, and the diagonal of $\mathbf{\Lambda}\in\mathbb{R^{\mathrm{m\times\mathrm{T}}}}$ consists of singular values.
Theoretical Justification.
According to Theorem 3.1, the bias vanishes as the correlations between labels to be aligned are eliminated. The obtained components are decorrelated (Lemma 3.2), thereby mitigating autocorrelation-induced bias. Moreover, component significance decreases from $\mathbf{Z}_{1}$ to $\mathbf{Z}_{\mathrm{T}}$ as they are derived by maximizing significance under sequentially augmented constraints. Furthermore, $\mathbf{P}^{*}$ can be computed via SVD (Lemma 3.3), offering an efficient alternative to sequentially solving the constrained optimization problems in (3).
Case study. To showcase the implications of the obtained components, a case study was conducted on the ETTh1 dataset. Implementation details are provided in Appendix A. The results are illustrated in 2(a), with key observations summarized as follows:
  • Decorrelation effect: 2(a) (a) compares the correlation volume in the label sequence and the generated components. In the left panel, the value at row i and column j represents the correlation between the i-th and j-th steps. A large number of non-diagonal elements exhibit substantial values, with approximately 50.5% exceeding 0.25, indicating notable autocorrelation in the label sequence. In contrast, the right panel shows negligible values for the non-diagonal elements. This demonstrates that transforming the label sequence into components effectively eliminates correlation, thereby corroborating Theorem 3.2.
  • Significance discrimination: 2(a) (b) compares the variance of the label sequence and associated components. In the left panel, the variance of different steps in the label sequence is relatively uniform, ranging from -1.5 to 1.5, suggesting that all steps are equally significant. In the right panel, however, only a few components exhibit large variance, while the others fluctuate within a narrow range. This indicates that the significance of different components can be clearly discerned, allowing for a trade-off between a slight loss of information and reduced optimization complexity by focusing on the most significant components.
The transformation is highly inspired by principal component analysis (PCA) (pca1, ; pca2, ). However, one key distinction warrants emphasis. Existing works dominantly employ principal component analysis on input features for obtaining informative representations (pca3, ; pca4, ), in contrast, we adapt it to label sequence, specifically aiming to reduce autocorrelation bias and simplify optimization for time-series forecasting. To our knowledge, this remains a technically innovative strategy.
Algorithm 1 The workflow of Time-o1.
Input: $\hat{\mathbf{Y}}$: forecast sequences, $\mathbf{Y}$: label sequences.
Parameter: $\alpha$: the relative weight of the transformed loss, $\gamma$: the ratio of involved significant components.
Output: $\mathcal{L}_{\alpha,\gamma}$: the obtained learning objective.
1:  $\mathbf{Y}\leftarrow\mathrm{Standardize}(\mathbf{Y})$.
2:  $\mathbf{P}^{*}\leftarrow\mathrm{SVD(\mathbf{Y})}$
3:  $\mathbf{Z}\leftarrow\mathbf{Y}\mathbf{P}^{*},\hat{\mathbf{Z}}\leftarrow\hat{\mathbf{Y}}\mathbf{P}^{*}$
4:  $\mathrm{K}\leftarrow\mathrm{round}(\gamma\cdot\mathrm{T})$
5:  $\mathcal{L}_{\mathrm{trans},\gamma}\leftarrow\|\hat{\mathbf{Z}}_{\cdot,1:\mathrm{K}}-\mathbf{Z}_{\cdot,1:\mathrm{K}}\|_{1}$
6:  $\mathcal{L}_{\mathrm{tmp}}\leftarrow\|\hat{\mathbf{Y}}-\mathbf{Y}\|_{2}^{2}$
7:  $\mathcal{L}_{\alpha,\gamma}:=\alpha\cdot\mathcal{L}_{\mathrm{trans},\gamma}+(1-\alpha)\cdot\mathcal{L}_{\mathrm{tmse}}.$
3.3 Model implementation
In this section, we present the implementation details of Time-o1. The approach centers on extracting the latent components from the label sequence, then optimizing the forecast model using the most significant components.
Given an input historical sequence, the forecast model predicts a sequence $\hat{\mathbf{Y}}$. In line with prevailing preprocessing practices itransformer ; DLinear ; fredformer , label sequences are first standardized (step 1), which facilitates the decorrelation prerequisite specified in Lemma 3.2. Next, following Lemma 3.3, we compute the optimal projection by applying SVD to the label sequence. The matrix $\mathbf{P}^{*}$, composed of the right singular vectors, provides the required projections described in (3). Both forecasted and label sequences are then projected into the latent component space (step 3), where the first column carries the largest significance, successively diminishing across columns.
Suppose $\mathrm{K}$ is the number of retained components, the training objective using them is given by:
$\mathcal{L}_{\mathrm{trans},\gamma}:$ $=\left\|\hat{\mathbf{Z}}_{\cdot,1:\mathrm{K}}-\mathbf{Z}_{\cdot,1:\mathrm{K}}\right\|_{1},$ (4)
where $\mathrm{K}=\mathrm{round}(\gamma\cdot\mathrm{T})$, with $\gamma$ controlling the involution ratio, the $\ell_{1}$ norm $\left\|\cdot\right\|_{1}$ computes the sum of element-wise absolute differences. Typically, we use the $\ell_{1}$ norm instead of the squared norm following (wang2025iclrfredf, ), considering that latent components typically vary greatly in scale (2(a)), which makes the squared norm unstable in practice. The $\ell_{1}$ norm yields more stable and robust optimization.
Finally, the two objectives ($\mathcal{L}_{\mathrm{tmse}}$ and $\mathcal{L}_{\mathrm{trans},\gamma}$) are fused following (wang2025iclrfredf, ), with $0\leq\alpha\leq 1$ controlling the relative contribution:
$\mathcal{L}_{\alpha,\gamma}:=\alpha\cdot\mathcal{L}_{\mathrm{trans},\gamma}+(1-\alpha)\cdot\mathcal{L}_{\mathrm{tmse}}.$ (5)
By projecting both forecasts and labels into decorrelated components, Time-o1 effectively reduces autocorrelation bias. By focusing exclusively on the most significant components, Time-o1 reduces optimization difficulty with minimal information loss. Time-o1 is model-agnostic, offering practitioners the flexibility to employ the most suitable forecast model for each specific scenario.
4 Experiments
To demonstrate the efficacy of Time-o1, there are six aspects empirically investigated:
  1. 1.
    Performance: Does Time-o1 work? We compare Time-o1 with state-of-the-art baselines using public datasets on long-term forecasting in Section 4.2 and short-term forecasting tasks in Appendix E.1. Moreover, we compare it with other learning objectives in Section 4.3.
  2. 2.
    Gain: How does it work? Section 4.4 offers an ablative study to dissect the contributions of the individual factors of Time-o1, elucidating their roles in enhancing forecast accuracy.
  3. 3.
    Generality: Does it support other forecast models? Section 4.5 verifies the adaptability of Time-o1 across different forecast models, with additional results in Appendix E.4.
  4. 4.
    Flexibility: Does it support alternative transformations? Section 4.5 also investigates generating latent components with other transformations to showcase flexibility of implementation.
  5. 5.
    Sensitivity: Does it require careful fine-tuning? Section 4.6 presents a sensitivity analysis of the hyperparameter $\alpha$, where Time-o1 maintains efficacy across a broad range of parameter values.
  6. 6.
    Efficiency: Is Time-o1 computationally expensive? Section D investigates the running cost of Time-o1 in diverse settings.
4.1 Setup
Datasets.
In this work, we conduct experiments on ETT (4 subsets), ECL, Traffic, Weather, and PEMS (itransformer, ) for long-term forecasting task, and M4 for short-term forecasting task Timesnet . All datasets are split chronologically into training, validation, and testing sets following the established work wang2025iclrfredf .
Baselines.
We compare Time-o1 against several established methods, grouped as: (1) Transformer-based methods: Transformer (Transformer, ), Autoformer (Autoformer, ), FEDformer (fedformer, ), iTransformer (itransformer, ), and Fredformer (fredformer, ); (2) MLP-based methods: DLinear (DLinear, ), TiDE (das2023long, ), and FreTS (FreTS, ); and (3) other competitive models: TimesNet (Timesnet, ) and MICN (micn, ).
Implementation.
The baseline models are reproduced using the scripts provided by Fredformer (fredformer, ). Notably, we disable the drop-last trick to ensure fair comparison following Qiu et al. qiu2024tfb . They are trained using the Adam (Adam, ) optimizer to minimize the TMSE loss. Datasets are split chronologically into training, validation, and test sets. Following the protocol outlined in the comprehensive benchmark (qiu2024tfb, ), the dropping-last trick is disabled during the test phase. When integrating Time-o1 to enhance an established model, we adhere to the associated hyperparameter settings in the public benchmark (fredformer, ; itransformer, ), only tuning $\alpha$, $\gamma$ and learning rate conservatively. Experiments are conducted on Intel(R) Xeon(R) Platinum 8383C CPUs and NVIDIA RTX H100 GPUs.
4.2 Overall performance
Table 1: Long-term forecasting performance.
Models
Time-o1
Fredformer
iTransformer
FreTS
TimesNet
MICN
TiDE
DLinear
FEDformer
Autoformer
Transformer
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
0.380
0.393
0.387 0.398
0.411
0.414
0.414
0.421
0.438
0.430
0.396
0.421
0.413
0.407
0.403
0.407
0.442
0.457
0.526
0.491
0.799
0.648
ETTm2
0.272
0.317
0.280 0.324
0.295
0.336
0.316
0.365
0.302
0.334
0.308
0.364
0.286
0.328
0.342
0.392
0.308
0.354
0.315
0.358
1.662
0.917
ETTh1
0.431
0.429
0.447 0.434
0.452
0.448
0.489
0.474
0.472
0.463
0.533
0.519
0.448
0.435
0.456
0.453
0.447
0.470
0.477
0.483
0.983
0.774
ETTh2
0.359
0.388
0.377
0.402
0.386
0.407
0.524
0.496
0.409
0.420
0.620
0.546
0.378
0.401
0.529
0.499
0.452
0.461
0.448
0.460
2.688
1.291
ECL
0.170
0.260
0.191
0.284
0.179 0.270
0.199
0.288
0.212
0.306
0.192
0.302
0.215
0.292
0.212
0.301
0.214
0.328
0.249
0.354
0.265
0.358
Traffic
0.419
0.280
0.486
0.336
0.426 0.285
0.538
0.330
0.631
0.338
0.529
0.312
0.624
0.373
0.625
0.384
0.640
0.398
0.662
0.416
0.692
0.379
Weather
0.241
0.280
0.261
0.282
0.269
0.289
0.249
0.293
0.271
0.295
0.264
0.321
0.272
0.291
0.265
0.317
0.326
0.372
0.319
0.365
0.699
0.601
PEMS03
0.097
0.208
0.146
0.260
0.122
0.233
0.149
0.261
0.126
0.230
0.106 0.223
0.316
0.370
0.216
0.322
0.152
0.275
0.411
0.475
0.122
0.226
PEMS08
0.141
0.237
0.171
0.271
0.149 0.247
0.174
0.275
0.152
0.243
0.153
0.258
0.318
0.378
0.249
0.332
0.226
0.312
0.422
0.456
0.240
0.261
  • Note: We fix the input length as 96 following (itransformer, ). Bold and underlined denote best and second-best results, respectively. Avg indicates average results over forecast horizons: T=96, 192, 336 and 720. Time-o1 employs the top-performing baseline on each dataset as its underlying forecast model.
Table 1 presents the long-term forecasting results. Time-o1 consistently improves base model performance. For example, on ETTh1, it reduces Fredformer’s MSE by 0.016. Similar gains across other datasets further validate its effectiveness. These results suggest that modifying the learning objective can yield improvements comparable to, or even exceeding, those from architectural advancements. We attribute this to two key aspects of Time-o1: its decorrelation effect for debiased training, and its discrimination on significant components, which simplifies optimization.
(a) The visualization of forecast sequence generated by DF and Time-o1. The left panels in (a) and (b) present label and forecast sequences, the right panels present the associated components.
Table 2: Comparable results with other objectives for time-series forecast.
Loss Time-o1 FreDF Koopman Dilate Soft-DTW DPTA DF
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Fredformer ETTm1 0.379 0.393 0.384 0.394 0.389 0.400 0.389 0.400 0.397 0.402 0.396 0.402 0.387 0.398
ETTh1 0.431 0.429 0.438 0.434 0.452 0.443 0.453 0.442 0.460 0.449 0.460 0.449 0.447 0.434
ECL 0.178 0.270 0.179 0.272 0.190 0.282 0.187 0.280 0.206 0.298 0.202 0.294 0.191 0.284
Weather 0.255 0.276 0.256 0.277 0.257 0.279 0.258 0.280 0.261 0.280 0.260 0.280 0.261 0.282
iTransformer ETTm1 0.395 0.401 0.405 0.405 0.413 0.416 0.407 0.412 0.417 0.415 0.416 0.415 0.411 0.414
ETTh1 0.438 0.434 0.442 0.437 0.455 0.451 0.452 0.448 0.470 0.457 0.463 0.454 0.452 0.448
ECL 0.170 0.260 0.176 0.264 0.178 0.269 0.178 0.269 0.175 0.266 0.177 0.267 0.179 0.270
Weather 0.251 0.272 0.257 0.276 0.289 0.313 0.286 0.309 0.292 0.316 0.291 0.313 0.269 0.289
  • Note: Bold and underlined denote best and second-best results, respectively. The reported results are averaged over forecast horizons: T=96, 192, 336 and 720.
Showcases.
We visualize the forecast sequences and the generated components to showcase the improvements of Time-o1 in forecast quality. A snapshot on ETTm2 with historical window $\mathrm{H}=96$ and forecast horizon $\mathrm{T}=336$ is depicted in 4(a). Although the model trained using canonical DF captures general trends, its forecast struggles with large variations (e.g., peaks within steps 100-400). This reflects its difficulty in modeling significant, high-variance components. In contrast, Time-o1, by explicitly discriminating and aligning these significant components, generates a forecast that accurately captures these large variations, including the peaks within steps 100-400.
4.3 Learning objective comparison
Table 2 compares Time-o1 against other time-series learning objectives: FreDF (wang2025iclrfredf, ), Koopman (koopman, ), Dilate (le2019shape, ), Soft-DTW (soft-dtw, ), and DPTA (dtw, ). For fair evaluation, we integrated their official implementations into both Fredformer and iTransformer.
Overall, shape alignment objectives (Dilate, Soft-DTW, DPTA) offer little performance gain over canonical DF (using TMSE loss), consistent with the findings in (le2019shape, ). This phenomenon is rationalized by the fact that they do not mitigate the label autocorrelation nor reduce task amounts for simplifying optimization. FreDF improves performance by partly addressing autocorrelation bias. However, as discussed in Section 3.1, FreDF does not fully eliminate this bias, nor does it distinguish significant components to simplify the optimization landscape. Time-o1 directly addresses these two limitations of FreDF, leading to its superior overall performance.
4.4 Ablation studies
Table 3: Ablation study results.
Model Decorrelation Reduction Data T=96 T=192 T=336 T=720 Avg
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
DF ETTm1 0.326 0.361 0.365 0.382 0.396 0.404 0.459 0.444 0.387 0.398
ETTh1 0.377 0.396 0.437 0.425 0.486 0.449 0.488 0.467 0.447 0.434
ECL 0.150 0.242 0.168 0.259 0.182 0.274 0.214 0.304 0.179 0.270
Weather 0.174 0.228 0.213 0.266 0.270 0.316 0.337 0.362 0.249 0.293
Time-o1 ETTm1 0.338 0.366 0.369 0.383 0.397 0.403 0.458 0.441 0.391 0.398
ETTh1 0.376 0.395 0.437 0.430 0.478 0.450 0.469 0.467 0.440 0.436
ECL 0.150 0.239 0.164 0.253 0.178 0.268 0.210 0.296 0.175 0.264
Weather 0.170 0.216 0.213 0.259 0.262 0.300 0.332 0.351 0.244 0.281
Time-o1 ETTm1 0.324 0.359 0.362 0.379 0.390 0.400 0.451 0.438 0.382 0.394
ETTh1 0.373 0.395 0.433 0.423 0.476 0.445 0.474 0.463 0.439 0.431
ECL 0.147 0.238 0.162 0.252 0.174 0.267 0.205 0.294 0.172 0.263
Weather 0.172 0.220 0.211 0.259 0.261 0.301 0.331 0.353 0.244 0.283
Time-o1 ETTm1 0.321 0.357 0.360 0.378 0.389 0.400 0.447 0.435 0.379 0.393
ETTh1 0.368 0.391 0.424 0.422 0.467 0.441 0.465 0.463 0.431 0.429
ECL 0.145 0.235 0.159 0.249 0.173 0.264 0.203 0.292 0.170 0.260
Weather 0.169 0.219 0.210 0.258 0.259 0.297 0.327 0.349 0.241 0.280
  • Note: Bold and underlined denote best and second-best results, respectively.
Table 3 presents an ablation study dissecting the contributions of critical factors in Time-o1: the decorrelation effect and the task reduction effect. The main findings are summarized as follows.
  • Time-o1 improves DF by reducing the number of tasks to optimize. To this end, it employs a randomized matrix as the projection matrix to generate components and aligns only a subset of the obtained components. The involution ratio $\gamma$ is finetuned on the validation set. It consistently improves over DF (e.g., -0.012 MAE on Weather). This demonstrates that reducing tasks with a minimal loss of label information can reduce optimization difficulty and improve performance.
  • Time-o1 improves DF by aligning decorrelated components. To this end, the objective is calculated in (5) with $\gamma=1$. It also outperforms DF, achieving the second-best results overall. This demonstrates aligning decorrelated label components to mitigate bias benefits forecast performance.
  • Time-o1 integrates both factors above by aligning the most significant decorrelated components, which achieves the best performance, demonstrating the synergistic effect of these two factors.
4.5 Generalization studies
In this section, we investigate the utility of Time-o1 with different transformation strategies and forecast models, to showcase the generality of Time-o1. In the bar-plots, the forecast errors are averaged over forecast lengths (96, 192, 336, 720), with error bars as 50% confidence intervals.
(a) Improvement of Time-o1 applied to different forecast models, shown with colored bars for means over forecast lengths (96, 192, 336, 720) and error bars for 50% confidence intervals.
Varying transformations.
Table 4: Varying transformations results.
ECL Weather
Transformation MSE $\Delta$ MAE $\Delta$ MSE $\Delta$ MAE $\Delta$
None 0.179 - 0.270 - 0.249 - 0.293 -
RPCA 0.171 4.31% $\downarrow$ 0.261 3.16% $\downarrow$ 0.244 1.78% $\downarrow$ 0.286 2.38% $\downarrow$
SVD 0.175 2.24% $\downarrow$ 0.264 2.18% $\downarrow$ 0.248 0.34% $\downarrow$ 0.290 0.93% $\downarrow$
FA 0.175 2.35% $\downarrow$ 0.265 1.82% $\downarrow$ 0.245 1.35% $\downarrow$ 0.287 1.97% $\downarrow$
[…]
  • Note: $\Delta$ refers to the relative error reduction compared to the baseline (None). Bold and underlined denote best and second-best results.
We select alternative approaches to transform the label sequence into latent components and report the forecast performance in Table 4. The selected transformation methods include robust principal component analysis (RPCA) (rpca, ), SVD (svd, ), and factor analysis (FA, ). Noting that the output of SVD yields components here, not a projection matrix as in Section 3.2. Implementation details are in Appendix C. Overall, all these transformation methods outperform canonical DF without transformation. However, the components obtained by these methods, including RPCA, cannot be guaranteed to be decorrelated. Consequently, autocorrelation bias may persist. In contrast, our approach ensures full decorrelation of the derived components (see Lemma 3.2), effectively addressing autocorrelation bias and leading to the best overall performance.
Varying forecast models.
We explore the versatility of Time-o1 in augmenting representative forecast models: Fredformer (fredformer, ), iTransformer (itransformer, ), FreTS (FreTS, ), and DLinear (DLinear, ). As illustrated in 8(a), Time-o1 improves forecast performance in all cases. For instance, on the Weather dataset, iTransformer and FreTS with Time-o1 achieve substantial reductions in MSE—up to 6.9% and 2.9%, respectively. Further evidence of Time-o1’s versatility can be found in Appendix E.4. These results confirm Time-o1’s potential as a plug-and-play strategy to enhance adiverse forecast models.
4.6 Hyperparameter sensitivity
Table 5: Hyperparameter results on $\alpha$.
ETTm1
ETTh2
Weather
$\alpha$
MSE MAE MSE MAE MSE MAE
0 0.3867 0.3979 0.3766 0.4019 0.2486 0.2930
0.3 0.3871 0.3983 0.3742 0.3982 0.2439 0.2851
0.5 0.3864 0.3976 0.3703 0.3964 0.2432 0.2833
0.7 0.3831 0.3959 0.3674 0.3943 0.2433 0.2849
1 0.3850 0.3933 0.3606 0.3890 0.2753 0.3209
  • Note: Bold and underlined denote best and second-best results.
Table 6: Hyperparameter results on $\gamma$.
ETTm1
ETTh2
Weather
$\gamma$
MSE MAE MSE MAE MSE MAE
0.1 0.3915 0.4002 0.3816 0.4029 0.2437 0.2845
0.3 0.3849 0.3964 0.3694 0.3961 0.2424 0.2825
0.5 0.3817 0.3943 0.3651 0.3923 0.2466 0.2877
0.7 0.3798 0.3930 0.3603 0.3886 0.2443 0.2861
1 0.3814 0.3940 0.3624 0.3903 0.2491 0.2924
  • Note: Bold and underlined denote best and second-best results.
In this section, we examine the impact of critical hyperparameters on the performance of Time-o1. The results are presented in Table 6 and Table 6. Additional trends across different datasets and forecast lengths are provided in Appendix E.5. The primary observations are summarized as follows:
  • The coefficient $\alpha$ determines the relative importance of the transformed objective in (5). When $\alpha$ is set to 1, Time-o1 exclusively uses the transformed objective. We observe that increasing $\alpha$ from 0 to 1 generally leads to improved forecasting accuracy, with the best results typically achieved when $\alpha$ is close to 1. The performance improvement is significant, e.g., MSE reduction on ETTh2 by 0.016, showcasing the utility of the transformed objective to improve forecast performance.
  • The coefficient $\gamma$ determines the ratio of involved components for alignment. When $\gamma$ is set to 1, Time-o1 aligns all obtained components for model training. The results demonstrate that setting $\gamma$ to 1, with all label information preserved, does not necessarily yield optimal performance. Instead, the best results are often obtained at $\gamma<1$, rendering some loss of label information. For instance, $\gamma=0.7$ yields the best results on ETTm1 and ETTh2, while $\gamma=0.3$ is optimal for the Weather dataset. The rationale is that focusing on aligning the most significant components can reduce the task amount, thereby simplifying optimization. Since the majority of the information is contained in the most significant components, the information loss is minimal. Collectively, these factors contribute to improved forecast performance.
5 Conclusion
In this study, we highlight the importance of designing effective objectives for time-series forecasting. Two critical challenges are formulated: label autocorrelation, which induces bias, and the excessive number of tasks, which impedes optimization. To address these challenges, we introduce a model-agnostic learning objective called Time-o1. This method transforms the label sequence into decorrelated components with discernible significance. Forecast models are trained to align the most significant components, which effectively mitigates label autocorrelation due to the decorrelation between components and reduces task amount by discarding non-significant components. Experiments demonstrate that Time-o1 improves the performance of forecast models across diverse datasets.
Limitations & future works. In this work, we investigate the challenges of label autocorrelation and excessive number of tasks in time-series forecasting. Nevertheless, these issues also manifest in areas such as speech generation, target recognition, and dense image prediction. Applying Time-o1 in these contexts is a promising avenue for future research. Additionally, historical sequence also exhibits autocorrelation and contains redundancy. Transforming inputs to derive decorrelated, compact representations could offer additional performance gains and also warrants investigation.
Acknowledgments
This project was supported by National Natural Science Foundation of China (62276004, 623B2002).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect our paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Our limitations are discussed in conclusions.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All proofs of theoretical results are involved in Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We detail our training and evaluation protocols in the experimental setting section.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We use open access data, and the code is provided in an anonymous link.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We detail our training and evaluation protocols in the experimental setting section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In Figure 3 which are crucial for evaluating the efficacy of TransDF, we report the error bars to make the results more rigorous and comprehensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the type of compute workers, memory in Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Since it is an algorithm-oriented research, there is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not involve these issues.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators or original owners of assets (e.g., code, data, models), used in the paper, are properly credited. The license and terms of use are explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not introduce new datasets and benchmarks in the paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We use open-access datasets and do not involve human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We use open-access datasets and do not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We do not involve LLMs as any important, original, or non-standard components in this work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A On the Implementation Details of Label Correlation Estimation
In this section, we introduce the motivation and implementation details of the label autocorrelation estimation techniques in 2(a). Measuring label autocorrelation $Y_{t}\rightarrow Y_{t^{\prime}}$ is indeed challenging due to the presence of confounding effect [33, 32, 17]. Specifically, the fork structure $Y_{t}\leftarrow L\rightarrow Y_{t^{\prime}}$ introduces spurious correlations between $Y_{t}$ and $Y_{t^{\prime}}$, thereby distorting the true strength of the label autocorrelation $Y_{t}\rightarrow Y_{t^{\prime}}$ of interest. This structural confounding undermines the validity of traditional measures such as Pearson correlation for quantifying label autocorrelation.
The previous work [37] involved the double machine learning (DML) method to estimate the ground-truth correlation while mitigating the influence of the fork structure. We adopt this in our experiments. DML is a statistical technique designed to estimate the causal effect of a treatment on an outcome while controlling for fork variables. Specifically, suppose we have a treatment variable $\mathcal{T}$, an outcome variable $\mathcal{Y}$, and a set of fork variables $\mathcal{X}$. The goal is to estimate the causal effect of $\mathcal{T}$ on $\mathcal{Y}$ while controlling for the influence of $\mathcal{X}$. To this end, DML first orthogonalizes both the treatment and outcome with respect to the fork variables. Two parametric models are employed to predict the treatment and outcome based on the fork variables. These predictions capture the impact of $\mathcal{X}$ on $\mathcal{Y}$ and $\mathcal{T}$. Subsequently, such impact of $\mathcal{X}$ is eliminated by calculating the residuals. Finally, the DML method regresses the outcome residuals on the treatment residuals, thereby measuring the causal effect of $\mathrm{T}$ on $\mathrm{Y}$ while removing the influence of the fork variables.
In our experiments, we measure label autocorrelation by treating the input sequence $L$ as the fork variable and different steps of the label sequences $Y_{t}$ and $Y_{t^{\prime}}$ as the treatment and outcome variables, respectively. Then, we estimate the treatment effect of $Y_{t}$ on $Y_{t^{\prime}}$ controlling $L$. Similarly, when measuring the correlation between different components, we use different components $Z_{k}$ and $Z_{k^{\prime}}$ as the treatment and outcome variables. Linear regression model is employed as the parametric model for both the treatment and outcome variables for efficiency, which is consistent to [37].
(a) The label autocorrelation in the original label sequence and the extracted components. The datasets are ETTh1, ETTh2, ETTm1, and Weather from left to right. The forecast length is set to 96.
To further complement the case study in 2(a), we analyzed the correlation matrices of the label sequences and the extracted components across multiple datasets, with the results presented in 10(a). The main observations are summarized as follows.
  • Panel (a) displays the correlation matrix of the label sequence, characterized by substantial non-diagonal elements, which highlight the strong autocorrelation among the labels. In contrast, panel (b) shows the correlation matrix of the extracted components, where the non-diagonal elements are nearly zero, indicating effective decorrelation.
  • Compared to the results reported in [37], where some obtained components remain correlated, the non-diagonal elements in panel (b) are fully eliminated. This difference arises because the Fourier transform in [37] achieves decorrelation only when the original label sequence is nearly infinitely long ($\mathrm{T}\textrightarrow\infty$), a condition that is not met in real-world applications with finite forecast horizons. This limitation stems from the predefined nature of the projection matrix, which lacks adaptation to the specific properties of the data. In contrast, our method ensures decorrelation by solving a constrained optimization problem, without relying on an infinitely long forecast horizon, thereby providing a more reliable approach for handling autocorrelation bias.
Appendix B Theoretical Justification
Theorem B.1 (Autocorrelation bias, Theorem 3.1 in the main text).
Given label sequence $Y$ where $\Sigma\in\mathbb{R}^{\mathrm{T}\times\mathrm{T}}$ denotes the step-wise correlation coefficient, the TMSE in (1) is biased compared to the negative log-likelihood of the label sequence, which is given by:
$\mathrm{Bias}=\left\|Y-\hat{Y}\right\|_{\Sigma^{-1}}^{2}-\left\|Y-\hat{Y}\right\|^{2}-\frac{1}{2}\log\left|\Sigma\right|.$ (6)
where $\|v\|_{\Sigma^{-1}}^{2}=v^{\top}\Sigma^{-1}v$. The bias vanishes if different steps in $Y$ are decorrelated. [^3] The pioneering work [37] identifies the bias under the first-order Markov assumption on the label sequence. This study generalizes this bias without the first-order Markov assumption.
Proof.
The proof follows our previous work [37] but relaxes the first-order Markov assumption.
Suppose the label sequence follows a multivariate normal distribution with mean vector $\mu=[\hat{Y}_{1},\hat{Y}_{2},\ldots,\hat{Y}_{\mathrm{T}}]$ and covariance matrix $\Sigma$, where the off-diagonal entries are $\Sigma_{ij}=\rho_{ij}\sigma^{2}$ for $i\neq j$. Here, $\rho_{ij}$ denotes the partial correlation between $Y_{i}$ and $Y_{j}$ given the input sequence $L$. The log-likelihood of the label sequence $Y$ can be expressed as:
[…]
Removing the constant terms unrelated to $\hat{Y}$, we obtain the practical negative log-likelihood (PNLL):
[…]
On the other hand, the TMSE loss can be expressed as:
[…]
where $I$ is the identity matrix. The difference between TMSE and PNLL can be expressed as:
[…]
which immediately vanishes if the label sequence is decorrelated, i.e., $\Sigma=I$. The proof is completed. ∎
Lemma B.2 (Lemma 3.3 in the main text).
The projection matrix $\mathbf{P}^{*}$ can be obtained via singular value decomposition (SVD): $\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}(\mathbf{P}^{*})^{\top}$, where $\mathbf{U}\in\mathbb{R^{\mathrm{m\times m}}}$ and $\mathbf{P}^{*}\in\mathbb{R^{\mathrm{T\times T}}}$ consist of singular vectors, and the diagonal of $\mathbf{\Lambda}\in\mathbb{R^{\mathrm{m\times\mathrm{T}}}}$ consists of singular values.
Proof.
We first consider the case with $p=1$, where the orthogonal constrains are not involved. The Lagrangian can be written as:
$\mathcal{L}(\mathbf{P}_{1},\lambda_{1})=\mathbf{P}_{1}^{\top}\mathbf{S}\mathbf{P}_{1}-\lambda_{1}(\mathbf{P}_{1}^{\top}\mathbf{P}_{1}-1),$ (7)
where $\lambda_{1}$ is the Lagrangian multiplier. According to the first-order condition, the derivative with respect to $\mathbf{P}_{1}$ should be zero:
$\frac{\partial\mathcal{L}}{\partial\mathbf{P}_{1}}\Big|_{\mathbf{P}_{1}=\mathbf{P}_{1}^{*}}=2\mathbf{S}\mathbf{P}_{1}^{*}-2\lambda_{1}\mathbf{P}_{1}^{*}=0,$ (8)
which immediately follows by $\mathbf{S}\mathbf{P}_{1}^{*}=\lambda_{1}\mathbf{P}_{1}^{*}$. Apparently, $\mathbf{P}_{1}^{*}$ is an eigenvector of $\mathbf{S}$, with corresponding eigenvalue $\lambda_{1}$. Moreover, $\mathbf{P}_{1}^{*}$ is the leading eigenvector of $\mathbf{S}$ associated with the largest eigenvalue, which follows from the maximization objective is $\mathbf{P}_{1}{{}^{*}\top}\mathbf{S}\mathbf{P}_{1}^{*}=\lambda_{1}$,
We further consider the case with $p\geq 2$, impose orthogonality to all previous projection vectors. Defining Lagrangian multipliers $\lambda_{p}$ and $\{\mu_{j}\}_{j=1}^{p-1}$, we write the Lagrangian as follow
$\mathcal{L}(\mathbf{P}_{p},\lambda_{p},\{\mu_{j}\})=\mathbf{P}_{p}^{\top}\mathbf{S}\mathbf{P}_{p}-\lambda_{p}(\mathbf{P}_{p}^{\top}\mathbf{P}_{p}-1)-\sum_{j=1}^{p-1}\mu_{j}\mathbf{P}_{p}^{\top}\mathbf{P}_{j}.$ (9)
According to the first-order condition, the derivative with respect to $\mathbf{P}_{p}$ should be zero:
$\frac{\partial\mathcal{L}}{\partial\mathbf{P}_{p}}\Big|_{\mathbf{P}_{p}=\mathbf{P}_{p}^{*}}=2\mathbf{S}\mathbf{P}_{p}^{*}-2\lambda_{p}\mathbf{P}_{p}^{*}-\sum_{j=1}^{p-1}\mu_{j}\mathbf{P}_{j}=0.$ (10)
To resolve $\{\mu_{j}\}$, we take the inner product of both sides using $\mathbf{P}_{k}^{*\top}$ with k=1,2,,p1k=1,2,...,p-1:
$\mathbf{P}_{k}^{*\top}\mathbf{S}\mathbf{P}_{p}^{*}-\lambda_{p}\mathbf{P}_{k}^{*\top}\mathbf{P}_{p}^{*}-\frac{1}{2}\sum_{j<p}\mu_{j}\mathbf{P}_{k}^{*\top}\mathbf{P}_{j}=0.$ (11)
Since $\mathbf{P}_{k}^{*}$ is previously obtained that satisfies $\mathbf{S}\mathbf{P}_{k}^{*}=\lambda_{k}\mathbf{P}_{k}^{*}$ and $\mathbf{P}_{k}^{*\top}\mathbf{P}_{k}^{*}=1$, we have $\mathbf{P}_{k}^{*\top}\mathbf{S}\mathbf{P}_{p}^{*}=\lambda_{k}$. Due to the orthogonal constraint, the current projection vector should be orthogonal to the previously derived ones, we have $\mathbf{P}_{k}^{*\top}\mathbf{P}_{p}^{*}=0$ and $\mathbf{P}_{k}^{*\top}\mathbf{P}_{j}=\delta_{k,j}$, where $\delta_{k,j}=1$ if $j=k$ and 0 otherwise. Putting together, the equation simplifies to: $\mu_{k}=0$.
Since $\mu_{k}=0$ holds for all k=1,2,,p1k=1,2,...,p-1, we have $\mu_{1}=\mu_{2}=...=\mu_{p-1}=0$. Plugging back to (10), the optimal condition becomes:
$\mathbf{S}\mathbf{P}_{p}^{*}=\lambda_{p}\mathbf{P}_{p}^{*},$ (12)
with the additional restriction that $\mathbf{P}_{p}^{*}$ is orthogonal to all previous directions. Therefore, $\mathbf{P}_{p}^{*}$ must be the eigenvector of $\mathbf{S}$ corresponding to the $p$-th largest eigenvalue. That is, $\mathbf{P}$ can be derived by performing eigenvector decomposition on $\mathbf{S}$.
Moving forward, consider SVD: $\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top}$, we can represent $\mathbf{S}$ as
$\mathbf{S}=\mathbf{Y}^{\top}\mathbf{Y}=\mathbf{V}\mathbf{\Lambda}^{\top}\mathbf{U}^{\top}\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top}=\mathbf{V}\mathbf{\Lambda}^{2}\mathbf{V}^{\top},$ (13)
which implies that each column of $\mathbf{V}$ is the eigenvector of $\mathbf{S}$, i.e., $\mathbf{V}=\mathbf{P}^{*}$. Therefore, the projection matrix $\mathbf{P}^{*}$ can be obtained by performing SVD on $\mathbf{Y}$ as: $\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top}$, where $\mathbf{V}$ is exactly the optimum projection matrix $\mathbf{P}^{*}$. The proof is therefore completed. ∎
Lemma B.3 (Decorrelated components, Lemma 3.2 in the main text).
Suppose $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$ contains normalized label sequences for $\mathrm{m}$ samples, $\mathbf{Z}=[\mathbf{Z}_{1},...,\mathbf{Z}_{\mathrm{T}}]$ are the obtained components; for any $p\neq p^{\prime}$, we have $\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=0$.
Proof.
For any two latent components $\mathbf{Z}_{p}$ and $\mathbf{Z}_{p^{\prime}}$ with $p\neq p^{\prime}$, we have:
$\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=(\mathbf{Y}\mathbf{P}_{p})^{\top}(\mathbf{Y}\mathbf{P}_{p^{\prime}})=\mathbf{P}_{p}^{\top}\mathbf{Y}^{\top}\mathbf{Y}\mathbf{P}_{p^{\prime}}$ (14)
According to (12), $\mathbf{P}_{p}$ and $\mathbf{P}_{p^{\prime}}$ are eigenvectors of $\mathbf{Y}^{\top}\mathbf{Y}$, we have
$\mathbf{Y}^{\top}\mathbf{Y}\mathbf{P}_{p}=\lambda_{p}\mathbf{P}_{p},\qquad\mathbf{Y}^{\top}\mathbf{Y}\mathbf{P}_{p^{\prime}}=\lambda_{p^{\prime}}\mathbf{P}_{p^{\prime}},$ (15)
which immediately follows by $\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=\lambda_{p^{\prime}}\mathbf{P}_{p}^{\top}\mathbf{P}_{p^{\prime}}$. Recalling that different projection bases are constrained to orthogonal, i.e., $\mathbf{P}_{p}^{\top}\mathbf{P}_{p^{\prime}}=0$ for $p\neq p^{\prime}$, we have
$\mathbf{Z}_{p}^{\top}\mathbf{Z}_{p^{\prime}}=0\qquad\text{for all }p\neq p^{\prime}.$ (16)
The proof is completed. ∎
Appendix C Generalized Orthogonalization Methods
In this section, we introduce alternative transform methods for obtaining latent components, each with distinct characteristics such as dimensionality reduction and noise isolation. We discuss their implications for transforming label sequences in time-series forecasting, with a comparative study detailed in Section 4.5.
RPCA.
The robust principal component analysis decomposes the data into a low-rank informative component and a sparse noise component, effectively separating structured signals from noise. Specifically, given $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$, it is achieved by solving:
$\min_{\mathbf{V},\mathbf{S}}\|\mathbf{V}\|_{*}+\lambda\|\mathbf{S}\|_{1},\quad\text{subject to}\quad\mathbf{Y}=\mathbf{V}+\mathbf{S},$ (17)
where $\|\cdot\|_{*}$ is the nuclear norm, $\|\cdot\|_{1}$ is the element-wise $\ell_{1}$ norm, and $\lambda$ is a regularization parameter. Afterwards, it performs the principal component analysis on the obtained informative component $\mathbf{V}$ to derive the projection matrix $\mathbf{P}$. The latent components are generated by $\mathbf{Z}=\mathbf{Y}\mathbf{P}$. While this approach enhances noise elimination, it does not guarantee decorrelation of the derived components, as the projection matrix $\mathbf{P}$ is derived from $\mathbf{V}$ instead of the original data matrix $\mathbf{Y}$.
SVD.
The singular value decomposition provides a method to decompose the matrix into different components. Given $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$, we have:
$\mathbf{Y}=\mathbf{U}\mathbf{\Lambda}\mathbf{V}^{\top},$ (18)
where $\mathbf{U}\in\mathbb{R}^{\mathrm{m}\times\mathrm{r}}$ and $\mathbf{V}\in\mathbb{R}^{\mathrm{T}\times\mathrm{R}}$ are singular vectors, $\mathbf{\Lambda}\in\mathbb{R}^{\mathrm{r\times r}}$ is diagonal with rank $\mathrm{r}$. The right singular vector is used as the projection matrix, and the latent components are generated by $\mathbf{Z}=\mathbf{Y}\mathbf{V}$. One key distinction here needs to be highlighted. Unlike the workflow in the main text (Algorithm 1), the label sequence is not normalized after window generation before computing SVD here, resulting in non-decorrelated components.
FA.
Factor analysis models the observed data as linear combinations of a small number of latent factors plus noise, capturing the covariance structure through these unobserved factors. Specifically, given mean-centered $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$, the model assumes:
$\mathbf{Y}=\mathbf{V}\mathbf{F}^{\top}+\mathbf{E},$ (19)
where $\mathbf{V}\in\mathbb{R}^{\mathrm{m}\times\mathrm{K}}$ is the factor loading matrix, $\mathrm{K}$ is the number of latent factors ($\mathrm{K}\ll\mathrm{m}$), $\mathbf{F}\in\mathbb{R}^{\mathrm{T}\times\mathrm{K}}$ contains the latent factor scores for each sample, and $\mathbf{E}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$ is the noise matrix. The standard assumption is that each factor $f_{i}\sim\mathcal{N}(0,\mathbf{I})$ and noise $\epsilon_{i}\sim\mathcal{N}(0,\Psi)$, where $\Psi$ is a diagonal covariance matrix. The loadings $\mathbf{V}$ and factor scores $\mathbf{F}$ are typically estimated via maximum likelihood. The latent components are given by the estimated factor scores, i.e., $\mathbf{Z}=\mathbf{Y}\Psi^{-1}\mathbf{F}(\mathbf{I}+\mathbf{F}^{\top}\Psi^{-1}\mathbf{F})^{-1}:=\mathbf{Y}\mathbf{P}$  [^4] Adapted from source code of sklearn: https://github.com/scikit-learn/scikit-learn/blob/98ed9dc73/sklearn/decomposition/. This approach captures the covariance structure of $\mathbf{Y}$ via a small number of factors, but does not necessarily guarantee uncorrelated or noise-isolated components.
Appendix D Complexity Analysis
In this section, we analyze the running cost of Time-o1. The core computation of Time-o1 involves (a) calculating the projection matrix $\mathbf{P}^{*}$ via SVD, and (b) performing transformation on both predicted and label sequences, followed by calculating their point-wise MAE loss. Given the target matrix $\mathbf{Y}\in\mathbb{R}^{\mathrm{m}\times\mathrm{T}}$, the SVD step decomposes $\mathbf{Y}$ with an established complexity of $\mathcal{O}(\mathrm{m}\mathrm{T}^{2})$ (assuming $\mathrm{m}\geq\mathrm{T}$). For the sequence transformation, each sample (row) in $\mathbf{Y}$ is multiplied by the projection matrix $\mathbf{P}^{*}\in\mathbb{R}^{\mathrm{T}\times\mathrm{T}}$, resulting in a total complexity of $\mathcal{O}(\mathrm{m}\mathrm{T}^{2})$. The computation of point-wise MAE loss across all samples and forecast steps is $\mathcal{O}(\mathrm{m}\mathrm{T})$, which is negligible compared to the complexity of previous steps. Thus, the overall complexity per batch is dominated by the SVD and projection operations, both scaling as $\mathcal{O}(\mathrm{m}\mathrm{T}^{2})$. The main findings from the empirical evaluations are as follows.
  • 10(b) (a) presents the computational cost for calculating the projection matrix. Overall, it increases linearly with the sample size and quadratically with the prediction length, which aligns with the theoretical complexity. Importantly, this operation is performed only once before training begins, rendering the associated overhead acceptable.
  • 10(b) (b) presents the computational cost for the sequence transformation. The cost increases quadratically with the prediction length, but remains below 2 ms. This cost is comparable to that of a linear projection. Furthermore, sequence transformation is not required during inference.
In conclusion, Time-o1 does not add complexity to model inference, and the additional complexity during the training stage is negligible.
(b) Running cost for projection matrix calculation (left panel with varying number of samples, right panel with varying prediction length) and sequence transformation (left panel for forward pass, right panel for backward pass, with average and shaded areas for 95% confidence intervals).
Appendix E More Experimental Results
Table 7: The comprehensive results on the long-term forecasting task.
Models
Time-o1
Fredformer
iTransformer
FreTS
TimesNet
MICN
TiDE
DLinear
FEDformer
Autoformer
Transformer
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
96
0.321
0.357
0.326
0.361
0.338
0.372
0.342
0.375
0.368
0.394
0.319
0.366
0.353
0.374
0.346
0.373
0.401
0.434
0.485
0.468
0.503
0.482
192
0.360
0.378
0.365
0.382
0.382
0.396
0.385
0.400
0.406
0.409
0.364
0.395
0.391
0.393
0.380
0.390
0.415
0.446
0.504
0.482
0.807
0.664
336
0.389
0.400
0.396
0.404
0.427
0.424
0.416
0.421
0.454
0.444
0.395
0.425
0.423
0.414
0.413
0.414
0.432
0.450
0.520
0.489
0.847
0.678
720
0.447
0.435
0.459 0.444
0.496
0.463
0.513
0.489
0.527
0.474
0.505
0.499
0.486
0.448
0.472
0.450
0.522
0.500
0.594
0.523
1.037
0.771
Avg
0.379
0.393
0.387 0.398
0.411
0.414
0.414
0.421
0.438
0.430
0.396
0.421
0.413
0.407
0.403
0.407
0.442
0.457
0.526
0.491
0.799
0.648
ETTm2
96
0.172
0.251
0.177 0.260
0.182
0.265
0.188
0.279
0.184
0.262
0.178
0.277
0.182
0.265
0.188
0.283
0.205
0.289
0.218
0.300
0.386
0.441
192
0.235
0.294
0.242 0.300
0.257
0.315
0.264
0.329
0.257
0.308
0.266
0.343
0.247
0.304
0.280
0.356
0.271
0.334
0.282
0.340
1.410
0.881
336
0.293
0.333
0.302
0.340
0.320
0.354
0.322
0.369
0.315
0.345
0.299
0.354
0.307
0.343
0.375
0.420
0.327
0.366
0.335
0.370
1.940
1.070
720
0.388
0.389
0.399 0.397
0.423
0.411
0.489
0.482
0.452
0.421
0.489
0.482
0.408
0.398
0.526
0.508
0.428
0.425
0.423
0.420
2.914
1.276
Avg
0.272
0.317
0.280 0.324
0.295
0.336
0.316
0.365
0.302
0.334
0.308
0.364
0.286
0.328
0.342
0.392
0.308
0.354
0.315
0.358
1.662
0.917
ETTh1
96
0.368
0.391
0.377
0.396
0.385
0.405
0.398
0.409
0.399
0.418
0.381
0.416
0.387
0.395
0.389
0.404
0.391
0.433
0.449
0.465
1.028
0.778
192
0.424
0.422
0.437
0.425
0.440
0.437
0.451
0.442
0.452
0.451
0.497
0.489
0.439
0.425
0.442
0.440
0.418
0.448
0.459
0.469
1.010
0.789
336
0.467
0.441
0.486
0.449
0.480
0.457
0.501
0.472
0.488
0.469
0.589
0.555
0.482
0.447
0.488
0.467
0.487
0.484
0.511
0.500
0.908
0.743
720
0.465
0.463
0.488
0.467
0.504
0.492
0.608
0.571
0.549
0.515
0.665
0.617
0.484
0.471
0.505
0.502
0.494
0.514
0.488
0.498
0.987
0.785
Avg
0.431
0.429
0.447 0.434
0.452
0.448
0.489
0.474
0.472
0.463
0.533
0.519
0.448
0.435
0.456
0.453
0.447
0.470
0.477
0.483
0.983
0.774
ETTh2
96
0.282
0.330
0.293
0.344
0.301
0.349
0.315
0.374
0.321
0.358
0.351
0.398
0.291 0.340
0.330
0.383
0.351
0.391
0.355
0.397
1.485
0.959
192
0.359
0.381
0.372 0.391
0.383
0.397
0.466
0.467
0.418
0.417
0.492
0.489
0.376
0.392
0.439
0.450
0.456
0.456
0.478
0.471
4.218
1.585
336
0.394
0.414
0.420
0.433
0.425
0.432
0.522
0.502
0.464
0.454
0.656
0.582
0.417 0.427
0.589
0.538
0.477
0.492
0.459
0.469
2.775
1.361
720
0.400
0.427
0.421 0.439
0.436
0.448
0.792
0.643
0.434
0.450
0.981
0.718
0.429
0.446
0.757
0.626
0.522
0.505
0.499
0.502
2.274
1.257
Avg
0.359
0.388
0.377
0.402
0.386
0.407
0.524
0.496
0.409
0.420
0.620
0.546
0.378
0.401
0.529
0.499
0.452
0.461
0.448
0.460
2.688
1.291
ECL
96
0.145
0.235
0.161
0.258
0.150 0.242
0.180
0.266
0.170
0.272
0.170
0.281
0.197
0.274
0.197
0.282
0.187
0.302
0.189
0.304
0.253
0.350
192
0.159
0.249
0.174
0.269
0.168 0.259
0.184
0.272
0.183
0.282
0.185
0.297
0.197
0.277
0.197
0.286
0.207
0.322
0.271
0.371
0.262
0.356
336
0.173
0.264
0.194
0.290
0.182 0.274
0.199
0.290
0.203
0.302
0.190
0.298
0.212
0.292
0.209
0.301
0.211
0.326
0.243
0.352
0.269
0.363
720
0.203
0.292
0.235
0.319
0.214 0.304
0.234
0.322
0.294
0.366
0.221
0.329
0.254
0.325
0.245
0.334
0.253
0.361
0.295
0.388
0.277
0.365
Avg
0.170
0.260
0.191
0.284
0.179 0.270
0.199
0.288
0.212
0.306
0.192
0.302
0.215
0.292
0.212
0.301
0.214
0.328
0.249
0.354
0.265
0.358
Traffic
96
0.393
0.265
0.461
0.327
0.397 0.271
0.531
0.323
0.590
0.316
0.498
0.298
0.646
0.386
0.649
0.397
0.588
0.367
0.575
0.356
0.689
0.396
192
0.410
0.275
0.470
0.326
0.416 0.279
0.519
0.321
0.624
0.336
0.521
0.309
0.599
0.362
0.598
0.371
0.613
0.377
0.647
0.394
0.710
0.388
336
0.421
0.280
0.492
0.338
0.429 0.286
0.529
0.327
0.641
0.345
0.529
0.314
0.606
0.363
0.605
0.373
0.640
0.398
0.694
0.446
0.687
0.366
720
0.451
0.298
0.521
0.353
0.462 0.303
0.573
0.346
0.670
0.356
0.567
0.326
0.643
0.383
0.646
0.395
0.718
0.450
0.731
0.468
0.681
0.366
Avg
0.419
0.280
0.486
0.336
0.426 0.285
0.538
0.330
0.631
0.338
0.529
0.312
0.624
0.373
0.625
0.384
0.640
0.398
0.662
0.416
0.692
0.379
Weather
96
0.169
0.219
0.180
0.220
0.171
0.210
0.174
0.228
0.183
0.229
0.179
0.244
0.192
0.232
0.194
0.253
0.235
0.310
0.233
0.306
0.423
0.448
192
0.210
0.258
0.222
0.258
0.246
0.278
0.213
0.266
0.242
0.276
0.242
0.310
0.240
0.270
0.238
0.296
0.295
0.353
0.286
0.347
0.664
0.585
336
0.259
0.297
0.283
0.301
0.296
0.313
0.270
0.316
0.293
0.312
0.273
0.330
0.292
0.307
0.282
0.332
0.364
0.397
0.346
0.385
0.848
0.686
720
0.327
0.349
0.358
0.348
0.362
0.353
0.337
0.362
0.366
0.361
0.360
0.399
0.364
0.353
0.347
0.385
0.411
0.429
0.412
0.420
0.861
0.685
Avg
0.241
0.280
0.261
0.282
0.269
0.289
0.249
0.293
0.271
0.295
0.264
0.321
0.272
0.291
0.265
0.317
0.326
0.372
0.319
0.365
0.699
0.601
PEMS03
12
0.070
0.176
0.081
0.191
0.072 0.179
0.085
0.198
0.094
0.201
0.096
0.217
0.117
0.226
0.105
0.220
0.108
0.229
0.233
0.366
0.106
0.206
24
0.087
0.198
0.121
0.240
0.104
0.217
0.129
0.244
0.116
0.221
0.095 0.210
0.233
0.322
0.183
0.297
0.131
0.255
0.405
0.485
0.117
0.221
36
0.105
0.219
0.180
0.292
0.137
0.251
0.173
0.286
0.134
0.237
0.107 0.223
0.379
0.418
0.258
0.361
0.159
0.285
0.327
0.415
0.127
0.233
48
0.124
0.238
0.201
0.316
0.174
0.285
0.207
0.315
0.161
0.262
0.125 0.242
0.535
0.516
0.319
0.410
0.209
0.331
0.679
0.634
0.139
0.245
Avg
0.097
0.208
0.146
0.260
0.122
0.233
0.149
0.261
0.126
0.230
0.106 0.223
0.316
0.370
0.216
0.322
0.152
0.275
0.411
0.475
0.122
0.226
PEMS08
12
0.081
0.183
0.091
0.199
0.084 0.187
0.096
0.205
0.111
0.208
0.161
0.274
0.121
0.233
0.113
0.225
0.163
0.258
0.232
0.334
0.204
0.232
24
0.117
0.218
0.138
0.245
0.123 0.227
0.151
0.258
0.139
0.232
0.127
0.237
0.232
0.325
0.199
0.302
0.197
0.288
0.545
0.550
0.232
0.251
36
0.157
0.253
0.199
0.303
0.170
0.268
0.203
0.303
0.168 0.260
0.148
0.252
0.376
0.427
0.295
0.371
0.241
0.326
0.379
0.436
0.246
0.263
48
0.207 0.294
0.255
0.338
0.218
0.306
0.247
0.334
0.189
0.272
0.175
0.270
0.543
0.527
0.389
0.429
0.302
0.375
0.531
0.502
0.278
0.297
Avg
0.141
0.237
0.171
0.271
0.149
0.247
0.174
0.275
0.152
0.243
0.153
0.258
0.318
0.378
0.249
0.332
0.226
0.312
0.422
0.456
0.240
0.261
$1^{\text{st}}$ Count
43
42
0
1
0
1
0
0
1 1 1
0
0
0
0
0
0
0
0
0
0
0
  • Note: We fix the input length as 96 following [21]. Bold typeface highlights the top performance for each metric, while underlined text denotes the second-best results. Avg indicates the results averaged over forecasting lengths: T=96, 192, 336 and 720.
Table 8: The comprehensive results on the short-term forecasting task.
Models
Time-o1
Fredformer
iTransformer
FreTS
MICN
DLinear
Fedformer
Metric
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
SMAPE
MASE
OWA
Yearly
13.485
3.010
0.791
13.509 3.028 0.794
13.797
3.143
0.818
13.576
3.068
0.801
14.594
3.392
0.873
14.307
3.094
0.827
13.648
3.089
0.806
Quarterly
10.105
1.180
0.889
10.140 1.185 0.893
10.503
1.248
0.932
10.361
1.223
0.916
11.417
1.385
1.023
10.500
1.237
0.928
10.612
1.246
0.936
Monthly
12.649
0.930
0.875
12.696 0.931 0.878
13.227
1.013
0.935
13.088
0.990
0.919
13.834
1.080
0.987
13.362
1.007
0.937
14.181
1.105
1.011
Others
4.852
3.274
1.027
4.848
3.230
1.019
5.101
3.419
1.076
5.563
3.71
1.17
6.137
4.201
1.308
5.12
3.649
1.114
4.823
3.243 1.019
Average
11.841
1.585
0.851
11.879 1.590 0.854
12.298
1.68
0.893
12.169
1.66
0.883
13.044
1.841
0.962
12.48
1.674
0.898
12.734
1.702
0.914
$1^{\text{st}}$ Count
4
4
4
0
1 1
0
0
0
0
0
0
0
0
0
0
0
0
1
0
0
  • Note: Bold typeface highlights the top performance for each metric, while underlined text denotes the second-best results. Avg indicates the results averaged over forecasting lengths: yearly, quarterly, and monthly.
E.1 Long-term forecast performance
Additional results on long-term forecast performance are available in Table 7.
E.2 Short-term forecast performance
Additional results on short-term forecast performance are available in Table 8, where Fredformer [25] serves as the forecast model.
E.3 Showcases
Additional results on showcases are available in 14(a) and 14(b).
(a) The forecast sequences generated with DF and Time-o1. The forecast length is set to 336 and the experiment is conducted on ETTm2.
(b) The forecast sequences generated with DF and Time-o1. The forecast length is set to 192 and the experiment is conducted on ECL.
E.4 Generalization studies
Additional results on varying forecast models and transformations are available in 14(c) and Table 9.
(c) Performance of different forecast models with and without Time-o1. The forecast errors are averaged over forecast lengths and the error bars represent 50% confidence intervals.
Table 9: Varying transformation results.
Trans
PCA
RPCA
SVD
FA
DF
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ECL
96
0.1449
0.2348
0.1450 0.2349
0.1450
0.2350
0.1478
0.2385
0.1500
0.2415
192
0.1592
0.2487
0.1594 0.2487
0.1595
0.2490
0.1619
0.2517
0.1681
0.2591
336
0.1731 0.2645
0.1732
0.2646
0.1730
0.2643
0.1789
0.2711
0.1823
0.2744
720
0.2033
0.2920
0.2066 0.2960
0.2214
0.3066
0.2095
0.2975
0.2145
0.3035
Avg
0.1701
0.2600
0.1710 0.2611
0.1747
0.2637
0.1745
0.2647
0.1787
0.2696
Weather
96
0.1692
0.2185
0.1715 0.2199
0.1723
0.2223
0.1717
0.2247
0.1737
0.2277
192
0.2102
0.2575
0.2116
0.2590 0.2116
0.2597
0.2125
0.2636
0.2128
0.2661
336
0.2586
0.2971
0.2631 0.3072
0.2676
0.3110
0.2613
0.2997
0.2705
0.3159
720
0.3271
0.3487
0.3303 0.3581
0.3394
0.3681
0.3354
0.3610
0.3372
0.3623
Avg
0.2413
0.2805
0.2441 0.2860
0.2477
0.2903
0.2452
0.2872
0.2486
0.2930
E.5 Hyperparameter sensitivity
Additional results on hyperparameter sensitivity are available in 20(a) for $\alpha$ and 20(b) for $\gamma$.
(a) Time-o1 improves Fredformer performance given a wide range of transformed loss strength $\alpha$. These experiments are conducted on ETTh1 (a), ETTh2 (b), ETTm1 (c), ETTm2 (d), ECL (e), Weather (f) datasets. Different columns correspond to different forecast lengths (from left to right: 96, 192, 336, 720, and their average with shaded areas being 15% confidence intervals).
(b) Time-o1 improves Fredformer performance given a wide range of rank ratio $\gamma$. These experiments are conducted on ETTh1 (a), ETTh2 (b), ETTm1 (c), ETTm2 (d), ECL (e), and Weather (f) datasets. Different columns correspond to different forecast lengths (from left to right: 96, 192, 336, 720, and their average with shaded areas being 15% confidence intervals).
E.6 Comparison with different learning objectives
Additional results on comparing different learning objectives are available in Table 10.
Table 10: Comparable results with different learning objectives.
Loss Time-o1 FreDF Koopman Dilate Soft-DTW DPTA DF
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Forecast model: FredFormer
ETTm1 96 0.321 0.357 0.326 0.355 0.335 0.368 0.337 0.367 0.332 0.363 0.332 0.364 0.326 0.361
192 0.360 0.378 0.363 0.380 0.366 0.384 0.364 0.384 0.370 0.386 0.370 0.386 0.365 0.382
336 0.389 0.400 0.392 0.400 0.399 0.408 0.397 0.406 0.406 0.409 0.409 0.410 0.396 0.404
720 0.447 0.435 0.455 0.440 0.456 0.441 0.457 0.443 0.478 0.450 0.476 0.448 0.459 0.444
Avg 0.379 0.393 0.384 0.394 0.389 0.400 0.389 0.400 0.397 0.402 0.396 0.402 0.387 0.398
ETTh1 96 0.368 0.391 0.370 0.392 0.375 0.397 0.378 0.399 0.376 0.398 0.378 0.399 0.377 0.396
192 0.424 0.422 0.436 0.437 0.438 0.434 0.439 0.435 0.439 0.435 0.438 0.433 0.437 0.425
336 0.467 0.441 0.473 0.443 0.473 0.455 0.481 0.453 0.484 0.455 0.486 0.455 0.486 0.449
720 0.465 0.463 0.474 0.466 0.523 0.487 0.516 0.482 0.542 0.510 0.538 0.510 0.488 0.467
Avg 0.431 0.429 0.438 0.434 0.452 0.443 0.453 0.442 0.460 0.449 0.460 0.449 0.447 0.434
ECL 96 0.151 0.245 0.152 0.247 0.166 0.263 0.158 0.253 0.168 0.266 0.158 0.253 0.161 0.258
192 0.166 0.256 0.166 0.257 0.174 0.267 0.170 0.263 0.218 0.313 0.216 0.307 0.174 0.269
336 0.181 0.274 0.183 0.278 0.188 0.280 0.190 0.286 0.197 0.291 0.199 0.295 0.194 0.290
720 0.213 0.304 0.216 0.304 0.232 0.318 0.229 0.316 0.240 0.322 0.235 0.322 0.235 0.319
Avg 0.178 0.270 0.179 0.272 0.190 0.282 0.187 0.280 0.206 0.298 0.202 0.294 0.191 0.284
Weather 96 0.171 0.208 0.174 0.213 0.174 0.214 0.173 0.214 0.173 0.213 0.179 0.219 0.180 0.220
192 0.219 0.253 0.219 0.254 0.220 0.256 0.225 0.260 0.220 0.255 0.223 0.257 0.222 0.258
336 0.277 0.295 0.278 0.296 0.280 0.298 0.280 0.299 0.281 0.296 0.281 0.298 0.283 0.301
720 0.353 0.346 0.354 0.347 0.354 0.347 0.355 0.348 0.369 0.355 0.356 0.347 0.358 0.348
Avg 0.255 0.276 0.256 0.277 0.257 0.279 0.258 0.280 0.261 0.280 0.260 0.280 0.261 0.282
Forecast model: itranFormer
ETTm1 96 0.323 0.358 0.334 0.365 0.350 0.382 0.342 0.376 0.339 0.373 0.341 0.375 0.338 0.372
192 0.371 0.388 0.381 0.390 0.389 0.400 0.381 0.396 0.383 0.395 0.383 0.395 0.382 0.396
336 0.408 0.407 0.417 0.412 0.425 0.423 0.418 0.418 0.429 0.423 0.429 0.423 0.427 0.424
720 0.477 0.450 0.489 0.453 0.489 0.458 0.487 0.457 0.516 0.469 0.512 0.467 0.496 0.463
Avg 0.395 0.401 0.405 0.405 0.413 0.416 0.407 0.412 0.417 0.415 0.416 0.415 0.411 0.414
ETTh1 96 0.378 0.393 0.378 0.395 0.392 0.411 0.385 0.405 0.387 0.405 0.386 0.405 0.385 0.405
192 0.428 0.423 0.428 0.423 0.446 0.442 0.440 0.437 0.443 0.439 0.441 0.439 0.440 0.437
336 0.473 0.450 0.470 0.447 0.483 0.461 0.480 0.457 0.494 0.464 0.489 0.462 0.480 0.457
720 0.473 0.469 0.490 0.484 0.501 0.491 0.504 0.492 0.557 0.520 0.538 0.509 0.504 0.492
Avg 0.438 0.434 0.442 0.437 0.455 0.451 0.452 0.448 0.470 0.457 0.463 0.454 0.452 0.448
ECL 96 0.145 0.235 0.149 0.238 0.151 0.243 0.150 0.241 0.149 0.241 0.149 0.240 0.150 0.242
192 0.159 0.249 0.163 0.251 0.167 0.257 0.168 0.259 0.164 0.255 0.166 0.257 0.168 0.259
336 0.173 0.264 0.179 0.268 0.182 0.275 0.181 0.274 0.180 0.274 0.180 0.272 0.182 0.274
720 0.203 0.292 0.212 0.297 0.212 0.300 0.212 0.300 0.207 0.296 0.212 0.300 0.214 0.304
Avg 0.170 0.260 0.176 0.264 0.178 0.269 0.178 0.269 0.175 0.266 0.177 0.267 0.179 0.270
Weather 96 0.163 0.202 0.170 0.208 0.206 0.257 0.208 0.259 0.207 0.252 0.209 0.258 0.171 0.210
192 0.214 0.248 0.219 0.252 0.264 0.300 0.252 0.285 0.264 0.303 0.258 0.291 0.246 0.278
336 0.274 0.294 0.279 0.296 0.309 0.326 0.311 0.328 0.314 0.333 0.312 0.331 0.296 0.313
720 0.351 0.344 0.358 0.347 0.377 0.369 0.374 0.364 0.384 0.377 0.383 0.373 0.362 0.353
Avg 0.251 0.272 0.257 0.276 0.289 0.313 0.286 0.309 0.292 0.316 0.291 0.313 0.269 0.289
E.7 Varying input length results
Table 11: Varying input sequence length results on the Weather dataset.
Models
Time-o1 iTransformer Time-o1 PatchTST
Metrics
MSE MAE MSE MAE MSE MAE MSE MAE
Input sequence length 96 96 0.163 0.202 0.171 0.210 0.175 0.213 0.200 0.244
192 0.214 0.248 0.246 0.278 0.224 0.257 0.229 0.263
336 0.274 0.294 0.296 0.313 0.276 0.296 0.287 0.303
720 0.351 0.344 0.362 0.353 0.353 0.346 0.363 0.353
Avg 0.250 0.272 0.269 0.289 0.257 0.278 0.270 0.291
192 96 0.163 0.205 0.168 0.215 0.158 0.199 0.164 0.208
192 0.210 0.248 0.213 0.253 0.204 0.242 0.225 0.269
336 0.259 0.287 0.265 0.294 0.257 0.286 0.287 0.308
720 0.334 0.338 0.341 0.345 0.332 0.337 0.341 0.345
Avg 0.241 0.270 0.247 0.277 0.238 0.266 0.254 0.283
336 96 0.157 0.203 0.162 0.213 0.150 0.196 0.156 0.206
192 0.199 0.246 0.211 0.256 0.196 0.241 0.222 0.277
336 0.251 0.287 0.260 0.295 0.246 0.282 0.251 0.285
720 0.324 0.338 0.332 0.341 0.320 0.333 0.327 0.338
Avg 0.233 0.268 0.241 0.276 0.228 0.263 0.239 0.277
720 96 0.161 0.213 0.172 0.225 0.152 0.201 0.154 0.207
192 0.205 0.250 0.220 0.268 0.198 0.248 0.205 0.254
336 0.254 0.292 0.282 0.311 0.248 0.284 0.248 0.288
720 0.318 0.339 0.337 0.351 0.313 0.335 0.317 0.339
Avg 0.235 0.274 0.253 0.289 0.228 0.267 0.231 0.272
Additional results on varying input lengths are available in Table 11—complementing the fixed length of 96 used in the main text.
E.8 Random seed sensitivity
Table 12: Experimental results ($\mathrm{mean}_{\pm\mathrm{std}}$) with varying seeds (2021-2025).
Dataset
ECL Weather
Models Time-o1 DF Time-o1 DF
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
96 0.145±0.000 0.235±0.000 0.150±0.001 0.242±0.001 0.164±0.001 0.203±0.001 0.190±0.012 0.232±0.014
192 0.160±0.001 0.249±0.001 0.166±0.002 0.257±0.002 0.216±0.002 0.250±0.001 0.240±0.011 0.272±0.010
336 0.174±0.002 0.266±0.002 0.181±0.001 0.273±0.001 0.274±0.001 0.294±0.001 0.293±0.003 0.310±0.003
720 0.205±0.001 0.293±0.001 0.216±0.004 0.303±0.003 0.353±0.002 0.344±0.001 0.361±0.002 0.352±0.001
Avg 0.171±0.001 0.261±0.001 0.178±0.001 0.269±0.001 0.252±0.001 0.273±0.001 0.271±0.003 0.292±0.003
Additional results on varying random seeds are available in Table 12.